### v3

# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

Abstract
The ProxSkip algorithm for distributed optimization is gaining increasing attention due to its proven benefits in accelerating communication complexity while maintaining robustness against data heterogeneity. However, existing analyses of ProxSkip are limited to the strongly convex setting and do not achieve linear speedup, where convergence performance increases linearly with respect to the number of nodes. So far, questions remain open about how ProxSkip behaves in the non-convex setting and whether linear speedup is achievable. In this paper, we revisit distributed ProxSkip and address both questions. We demonstrate that the leading communication complexity of ProxSkip is $\mathcal{O}(\nicefrac{{p\sigma^{2}}}{{n\epsilon^{2}}})$ for non-convex and convex settings, and $\mathcal{O}(\nicefrac{{p\sigma^{2}}}{{n\epsilon}})$ for the strongly convex setting, where $n$ represents the number of nodes, $p$ denotes the probability of communication, $\sigma^{2}$ signifies the level of stochastic noise, and $\epsilon$ denotes the desired accuracy level. This result illustrates that ProxSkip achieves linear speedup and can asymptotically reduce communication overhead proportional to the probability of communication. Additionally, for the strongly convex setting, we further prove that ProxSkip can achieve linear speedup with network-independent stepsizes.
Index Terms:
Distributed optimization, stochastic non-convex optimization, local update, linear speedup.
I Introduction
In this work, we consider the following distributed stochastic optimization problem by a group of agents $[n]:=\{1,2,\ldots,n\}$ connected over a network:
$f^{\star}=\min_{{\bf{x}}\in\mathbb{R}^{d}}\Big{[}f({\bf{x}}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}({\bf{x}})\Big{]},$ (1)
with $\text{with }f_{i}({\bf{x}})=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F_{i}({\bf{x}},\xi_{i})],$
where $\{\mathcal{D}_{i}\}^{n}_{i=1}$ represent data distributions, which can be heterogeneous across $n$ nodes, $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth local function accessed by node $i$. In this setup, a network of nodes (also referred to as agents, workers, or clients) collaboratively seeks to minimize the average of the nodes’ objectives. Solving problem (1) in a distributed manner has garnered considerable attention in recent years [1, 2, 3, 4, 5, 6]. These methods do not rely on a central coordinator and that communicate only with neighbors in an arbitrary communication topology. Nevertheless, distributed optimization algorithms may still face challenges arising from communication bottlenecks.
To reduce communication costs, many techniques have been proposed. These techniques include compressing models and gradients [7, 8, 9], using accelerated scheme [10, 11, 12, 13, 14, 15, 16], and implementing local updates [17, 18, 19, 20]. By applying these strategies, it is possible to reduce the amount of information exchanged between different nodes during training, thereby improving the efficiency of distributed training setups.
In this work, we mainly focus on performing local updates as means to reduce communication frequency. In centralized settings, local-SGD/FedAvg [19, 21, 22] has emerged as one of the most widely adopted optimization methods that employ local updates. However, when dealing with heterogeneous data, Local-SGD/FedAvg encounters the challenge of “client-drift.” This phenomenon arises from the diversity of functions on each node, causing each client to converge towards the minima of its respective function $f_{i}$, which may be significantly distant from the global optimum $f^{\star}$. To tackle this issue, several algorithms have been proposed, including Scaffold [23], Scaffold with momentum [24], FedLin [25], FedPD [26], FedDyn [27], VRL-SGD [28], FedGATE [29], SCALLION/SCAFCOM [30]. In distributed settings, local-DSGD has been introduced in [31]. Similarly to local-SGD, it also encounters the issue of client-drift when dealing with heterogeneous data. To mitigate the drift in Local-DSGD, several algorithms have been proposed. Notably, gradient-tracking (GT) based approaches, such as local-GT [36] and $K$-GT [37], have been developed. Additionally, algorithms based on ED/NIDS/D2 [38, 39, 40, 41], such as LED [42], have been introduced. Distinct from these periodic local updates methods [23, 25, 26, 31, 36, 37, 42], methods incorporating probabilistic local update are proposed such ProxSkip [43] and its extended versions, such as TAMUNA [44], CompressedScaffnew [45], LoCoDL [46], VR-ProxSkip [47], ODEProx [48], and RandProx [49].
It is known that ProxSkip does not depend on the heterogeneity of the data and exhibit linear convergence on distributed strongly convex problems in the absence of stochastic noise [43]. When the network is sufficiently well-connected, ProxSkip [43] and its extensions [44, 45, 46, 47, 48, 49] are gaining increasing attention due to their proven benefits in accelerating communication complexity. When deploying ProxSkip within the context of distributed learning, it becomes imperative to comprehend its behavior on non-convex tasks and its susceptibility to stochastic noise. However, existing ProxSkip convergence analyses focus on convex settings, and the main limitation of the existing analyses is the inability to prove that linear speedup in terms of the number of nodes. Notice that although [48] presents the ODEProx algorithm and gives a tighter analysis of ProxSkip in the strongly convex setting, this new analysis shares the same limitation as the original ProxSkip analysis, namely, the inability to achieve linear speedup. Achieving linear speedup is highly desirable for a distributed learning and optimization algorithm as it enables effective utilization of the massive parallelism inherent in large distributed stochastic optimization systems. Consequently, two fundamental open questions emerge:
(1) How does ProxSkip behave on non-convex tasks?
(2) Can we establish a linear speedup bound for ProxSkip in the presence of stochastic noise?
TABLE I: Comparisons with existing convergence rates of ProxSkip for distributed optimization. $\sigma^{2}$ is the variance of the stochastic gradient. $\zeta_{0}:=\max\{1-\alpha\mu,1-p^{2}\}$, the definition of $\zeta_{{\mathrm{new}}}$ and $\zeta$ can be found in [48] and Theorem 1, respectively. CVX, N-CVX, and S-CVX mean convex, non-convex, and strongly convex, respectively.
Reference convergence rate decentralized linear speedup
N-CVX CVX S-CVX
[44, 45, 46] no results no results $\zeta_{0}^{T}$, $\sigma^{2}=0$
[43, 47] no results no results $\zeta_{0}^{T}+\mathcal{O}(\alpha\sigma^{2})$
[48] no results no results $\zeta_{{\mathrm{new}}}^{T}+\mathcal{O}(\alpha\sigma^{2})$
[43] no results no results $\zeta^{T}$, $\sigma^{2}=0$
[49] no results $\mathcal{O}\left(\frac{1}{\alpha T}\right)$ $\zeta^{T}$, $\sigma^{2}=0$
this work $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right)$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right)$ $\mathcal{O}\left(\zeta^{T}+\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right)$
In this paper, we revisit ProxSkip for distributed stochastic optimization and provide answers to both questions. Specifically, we develop a new analysis with a novel proof technique under non-convex, convex, and strongly convex settings. Through this analysis, we obtain several new results that are comparable to the bounds of state-of-the-art distributed algorithms while achieving linear speedup bounds. We highlight our contributions as follows:
  • We establish the non-asymptotic convergence rate under stochastic non-convex, convex, and strongly convex settings of ProxSkip for problem (1). In particular, we prove that ProxSkip at iteration $T$ converges with rate
    (N-)CVX: $\mathcal{O}\bigg{(}\frac{1}{\alpha T}+\frac{\alpha^{2}}{(1-\lambda_{2})^{2}T}+\frac{\alpha\sigma^{2}}{n}+\frac{\sigma^{2}\alpha^{2}}{1-\lambda_{2}}\bigg{)},$
    S-CVX: $\mathcal{O}\bigg{(}(1-\alpha\mu)^{T}a_{0}+\frac{\alpha\sigma^{2}}{\mu n}+\frac{\sigma^{2}\alpha^{2}}{\mu(1-\lambda_{2})}\bigg{)},$
    where $\alpha$ is the stepsize of ProxSkip, $\sigma^{2}$ denotes the variance of the stochastic gradient, $1-\lambda_{2}$ is a topology-dependent quantity that approaches 00 for a large and sparse network, $\mu$ is the strongly convex constant, and $a_{0}$ is a constant that depends on the initialization. To the best of our knowledge, it is the first work that establishes the convergence rate of probabilistic distributed methods for non-convex settings. We offer a comparison of convergence rates of ProxSkip for (1) in Table I.
  • We prove that, after enough transient time, the expected communication complexity of ProxSkip is $\mathcal{O}(\nicefrac{{p\sigma^{2}}}{{n\epsilon^{2}}})$ (or $\tilde{\mathcal{O}}(\nicefrac{{p\sigma^{2}}}{{n\epsilon}})$ for S-CVX), where $\epsilon$ denotes the desired accuracy level, demonstrating that ProxSkip achieves linear speedup with respect to the number of nodes $n$. In addition, for the strongly convex setting, we further prove that ProxSkip can achieve linear speedup with network-independent stepsizes. The proposed new proof technique overcomes the analytical limitations of [43, 44, 45, 46, 47, 48, 49]. To the best of our knowledge, we prove for the first time that ProxSkip can achieve linear speedup.
  • We elucidate the effects of noise, local steps, and data heterogeneity on the convergence of ProxSkip in stochastic non-convex, convex, and strongly convex settings. We demonstrate the robustness of ProxSkip against data heterogeneity while enhancing communication efficiency by local updates. Furthermore, we show that the convergence rates exhibited by ProxSkip in stochastic settings are comparable with those of existing state-of-the-art distributed algorithms incorporating local updates [31, 35, 36, 37, 42] (see Table II).
II Setup
All vectors are column vectors unless otherwise stated. We let $\mathbf{x}_{i}^{t}\in\mathbb{R}^{d}$ represent the local state of node $i$ at the $t$-th iteration. For the sake of convenience in notation, we use bold capital letters to denote stacked variables. For instance,
$\mathbf{X}^{t}:=$ $\ [\mathbf{x}_{1}^{t},\mathbf{x}_{2}^{t},\ldots,\mathbf{x}_{n}^{t}]^{\sf T}\in\mathbb{R}^{n\times d},$
$\mathbf{G}^{t}:=$ $\ [\mathbf{g}_{1}^{t},\mathbf{g}_{2}^{t},\ldots,\mathbf{g}_{n}^{t}]^{\sf T}\in\mathbb{R}^{n\times d},$
$\nabla F(\mathbf{X}^{t}):=$ $\ [\nabla f_{1}(\mathbf{x}^{t}_{1}),\nabla f_{2}(\mathbf{x}^{t}_{2}),\ldots,\nabla f_{n}(\mathbf{x}^{t}_{n})]^{\sf T}\in\mathbb{R}^{n\times d}.$
II-A Network graph
In this work, we focus on distributed scenarios (undirected and connected network), where a network of $n$ nodes is interconnected by a graph with a set of edges $\mathcal{E}$, where node $i$ is connected to node $j$ if $(i,j)\in\mathcal{E}$. To describe the algorithm, we introduce the global mixing matrix $\mathbf{W}=[W_{ij}]$, where $W_{ij}=W_{ji}=0$ if $(i,j)\notin\mathcal{E}$, and $W_{ij}>0$ otherwise. We impose the following standard assumption on the mixing matrix.
Assumption 1.
The mixing matrix $\mathbf{W}\in[0,1]^{n\times n}$ is symmetric, doubly stochastic, and primitive. Let $\lambda_{1}=1$ denote the largest eigenvalue of the mixing matrix $\mathbf{W}$, and the remaining eigenvalues are denoted as $11>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{n}>-$ 1.
We introduce two quantities as follows:
$\mathbf{W}_{a}=\mathbf{I}-\nicefrac{{1}}{{2\chi}}(\mathbf{I}-\mathbf{W}),\ \mathbf{W}_{b}=(\mathbf{I}-\mathbf{W})^{1/2},$
where $\chi\geq 1$. Under Assumption 1, the matrix $\mathbf{W}_{a}$ is positive semi-definite and doubly stochastic. Furthermore, we have $\mathbf{I}-\mathbf{W}_{a}=\nicefrac{{1}}{{2\chi}}\mathbf{W}_{b}^{2}$, and $\mathbf{W}_{a}$ is well-conditioned when $\chi$ is large.
II-B Algorithm description
Recalling [43], the ProxSkip algorithm for the distributed stochastic optimization problem (1) can be written as
$\hat{\mathbf{Z}}^{t}$ $=\mathbf{X}^{t}-\alpha\mathbf{G}^{t}-\mathbf{Y}^{t},$ (2a)
$\mathbf{X}^{t+1}$ $=(1-\theta_{t})\hat{\mathbf{Z}}^{t}+\theta_{t}\mathbf{W}_{a}\hat{\mathbf{Z}}^{t},$ (2b)
$\mathbf{Y}^{t+1}$ $=\mathbf{Y}^{t}+\beta(\hat{\mathbf{Z}}^{t}-\mathbf{X}^{t+1}).$ (2c)
Here, $\alpha>0$ is the stepsize, $\beta>0$, $\mathbf{G}^{t}=[\mathbf{g}_{1}^{t},\mathbf{g}_{2}^{t},\ldots,\mathbf{g}_{n}^{t}]^{\sf T}\in\mathbb{R}^{n\times d}$ with $\mathbf{g}_{i}^{t}$ representing the stochastic gradient of $\nabla f_{i}(\mathbf{x}_{i}^{t})$, $\theta_{t}=1$ with probability $p$ and $\theta_{t}=0$ with probability $1-p$, and $\mathbf{Y}^{t}$ is the control variate. At each iteration $t\geq 0$, communication takes place with a probability $p\in(0,1]$. In the absence of communication, the update $\mathbf{X}^{t+1}=\mathbf{X}^{t}-\alpha\mathbf{G}^{t}-\mathbf{Y}^{t}$ is performed, while $\mathbf{Y}^{t+1}$ remains unchanged. This allows for multiple iterations of local computations to be performed between communication rounds. Decomposing the updates for individual nodes, we provide a detailed implementation in Algorithm 1.
Algorithm 1 ProxSkip for distributed stochastic optimization
1:  Input $\alpha>0$, $\beta>0$, $0<p\leq 1$, $\chi\geq 1$, initial iterates ${\bf{x}}^{0}_{i}={\bf{x}}^{0}\in\mathbb{R}^{d},~{}i=1,\dots,n$, initial dual variables ${\bf{y}}^{0}_{i}=0,~{}i=1,\dots,n$, weights for averaging ${\bf{W}}_{a}={\bf{I}}-\nicefrac{{1}}{{2\chi}}({\bf{I-W}}):=(\widetilde{W}_{ij})^{n}_{i,j=1}$.
2:  Flip coins $[\theta_{0},\ldots,\theta_{T-1}]$, where $\theta_{t}\in\{0,1\}$, with $\mathop{{\bf{P}}}(\theta_{t}=1)=p$.
3:  for $t=0,1,\dotsc,T-1$ every node do
4:     Sample $\xi_{i}^{t}$, compute gradient ${\bf{g}}_{i}^{t}=\nabla F_{i}({\bf{x}}^{t}_{i},\xi^{t}_{i})$
5:     $\hat{{\bf{z}}}^{t}_{i}={\bf{x}}^{t}_{i}-\alpha{\bf{g}}^{t}_{i}-{\bf{y}}^{t}_{i}$triangleright update the prediction variate $\hat{{\bf{z}}}^{t}_{i}$
6:     if $\theta_{t}=1$ then
7:        ${\bf{x}}^{t+1}_{i}=\sum_{j=1}^{n}\widetilde{W}_{ij}\hat{{\bf{z}}}^{t}_{j}$triangleright communicate with probability $p$
8:        ${\bf{y}}^{t+1}_{i}={\bf{y}}^{t}_{i}+\beta(\hat{{\bf{z}}}^{t}_{i}-{\bf{x}}^{t+1}_{i})$triangleright update the control variate ${\bf{y}}^{t+1}_{i}$
9:     else
10:        ${\bf{y}}^{t+1}_{i}={\bf{y}}^{t}_{i},~{}{\bf{x}}^{t+1}_{i}=\hat{{\bf{z}}}^{t}_{i}$triangleright skip communication
11:     end if
12:  end for
II-C Assumptions
We further use the following standard assumptions:
Assumption 2.
A solution exists to problem (1), and $f^{*}>-\infty$. Moreover, $f_{i}$ is $L$-smooth, i.e.,
$\|\nabla f_{i}({\bf{x}})-\nabla f_{i}({\bf{y}})\|\leq L\|{\bf{x}}-{\bf{y}}\|,\forall{\bf{x}},{\bf{y}}\in\mathbb{R}^{d}.$
Assumption 3.
Each function $f_{i}$ is $\mu$-strongly convex for constant $\mu\geq 0$, i.e.,
$f_{i}({\bf{x}})-f_{i}({\bf{y}})+\frac{\mu}{2}\|{\bf{x}}-{\bf{y}}\|^{2}\leq\langle\nabla f_{i}({\bf{x}}),{\bf{x}}-{\bf{y}}\rangle,\forall{\bf{x}},{\bf{y}}\in\mathbb{R}^{d}.$
Assumption 4.
For all iteration $t\geq 0$, the local stochastic gradient ${\bf{g}}^{t}_{i}=\nabla F_{i}({\bf{x}}_{i}^{t},\xi^{t}_{i})$ is an unbiased estimate, i.e.,
$\mathbb{E}_{\xi^{t}_{i}}[\nabla F_{i}({\bf{x}}_{i}^{t},\xi^{t}_{i})\;|\;{\bf{x}}_{i}^{t}\ ]=\nabla f_{i}({\bf{x}}_{i}^{t}),$
and there exists a constant $\sigma>0$ such that
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi_{i}^{t}}\big{[}\|\nabla F_{i}({\bf{x}}_{i}^{t},\xi^{t}_{i})-\nabla f_{i}({\bf{x}}_{i}^{t})\|^{2}\big{]}\leq\sigma^{2}.$
III Convergence results
We now present our novel convergence results for ProxSkip. In Section III-A, we recall the existing results in [43]. In Section III-B, the convergence rates and communication complexities for nonconvex and convex functions are presented Theorem 1 and Corollary 1, respectively. In Section III-C, we prove further that ProxSkip can achieve linear speedup with network-independent stepsizes.
III-A Preliminary
We start to recall the existing convergence results of ProxSkip [43, 49].
Lemma 1.
Suppose that Assumptions 1, 2, and 4 hold, and $f_{i}$ is $\mu$-strongly convex for some $0<\mu\leq L$. If $0<\alpha\leq\nicefrac{{1}}{{L}}$, $\beta=p$, and $\chi\geq 1$, it holds that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq\zeta^{t+1}a_{0}+\frac{\alpha^{2}\sigma^{2}}{1-\zeta},$ (3)
where $a_{0}$ is a constant that depends on the initialization and $\zeta=\max\left\{1-\alpha\mu,1-\frac{(1-\lambda_{2})p^{2}}{2\chi}\right\}<1$.
When $\sigma^{2}=0$, by setting $\alpha=\nicefrac{{1}}{{L}}$ and $\chi=1$, we can deduce from (3) that the communication complexity of ProxSkip to achieve $\epsilon$-accuracy, i.e., $\mathbb{E}[|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}]\leq\epsilon$, is given by $\mathcal{O}((p\kappa+\nicefrac{{1}}{{p(1-\lambda_{2})}})\mathrm{log}\nicefrac{{1}}{{\epsilon}})$, where $\kappa=\nicefrac{{L}}{{\mu}}$. If the network is sufficiently well-connected, i.e., $\nicefrac{{1}}{{(1-\lambda_{2})\kappa}}<1$, and we set $p=\sqrt{\nicefrac{{1}}{{(1-\lambda_{2})\kappa}}}$, the iteration complexity becomes $\mathcal{O}(\sqrt{\nicefrac{{\kappa}}{{1-\lambda_{2}}}}\ \mathrm{log}\nicefrac{{1}}{{\epsilon}})$, achieving the optimal communication complexity as proven by [50].
When $\sigma^{2}\neq 0$, based on (3) and the fact that $\frac{\alpha^{2}\sigma^{2}}{1-\zeta}=\frac{\alpha^{2}\sigma^{2}}{\alpha\mu}=\mathcal{O}(\alpha\sigma^{2})$, we can conclude that the local solution ${\bf{x}}_{i}^{t}$ generated by ProxSkip converges to the global minimizer ${\bf{x}}^{\star}$ at a linear rate until it reaches an $\mathcal{O}(\alpha\sigma^{2})$-neighborhood of ${\bf{x}}^{\star}$. However, it is important to note that relying solely on equation (3) is not sufficient to achieve the desired linear speedup term $\mathcal{O}(\frac{\alpha\sigma^{2}}{n})+\mathcal{O}(\alpha^{2})$. This indicates that the direct extension of the analysis techniques proposed in [43] or [49] to the stochastic scenario does not guarantee linear speedup, despite ensuring convergence. Therefore, further analysis is required to achieve the desired linear speedup.
III-B Main theorem—Convergence rate of ProxSkip
We are now ready to present the new convergence results for ProxSkip.
Theorem 1.
Suppose that Assumptions 1, 2, and 4 hold. Let $\bar{{\bf{x}}}^{t}=\frac{1}{n}\sum_{i=1}^{n}{\bf{x}}_{i}^{t}$ denote the iterates of Algorithm 1 and ${\bf{x}}^{\star}$ solves (1). For sufficiently small $\alpha$, $\chi=\mathcal{O}(\max\{1,\nicefrac{{(1-p)}}{{(1-\lambda_{2})}}\})$, and $\beta=1$, we have the following convergence results.
Non-convex: Let $F_{0}=f(\bar{{\bf{x}}}^{0})-f^{\star}$ and $\varsigma^{2}_{0}=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\bar{{\bf{x}}}^{0})-\nabla f(\bar{{\bf{x}}}^{0})\|^{2}$. It holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]$
$\leq\mathcal{O}\bigg{(}\underbrace{\frac{F_{0}}{\alpha T}+\frac{\chi^{2}L^{2}\varsigma^{2}_{0}\alpha^{2}}{(1-\lambda_{2})^{2}T}}_{{\mathrm{deterministic\ part}}}+\underbrace{\frac{\alpha\sigma^{2}L}{n}+\frac{\chi L^{2}\sigma^{2}\alpha^{2}}{1-\lambda_{2}}}_{{\mathrm{stochastic\ part}}}\bigg{)}.$ (4)
Convex: Let $R_{0}^{2}=\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}$. Under the additional Assumption 3 with $\mu\geq 0$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]$
$\leq\mathcal{O}\bigg{(}\underbrace{\frac{R_{0}^{2}}{\alpha T}+\frac{\chi^{2}L\varsigma^{2}_{0}\alpha^{2}}{(1-\lambda_{2})^{2}T}}_{{\mathrm{deterministic\ part}}}+\underbrace{\frac{\alpha\sigma^{2}}{n}+\frac{\chi L\sigma^{2}\alpha^{2}}{1-\lambda_{2}}}_{{\mathrm{stochastic\ part}}}\bigg{)}.$ (5)
Strongly convex: Under the additional Assumption 3 with $\mu>0$, it holds that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq$ $\underbrace{\Big{(}1-\frac{\alpha\mu}{4}\Big{)}^{T}\Big{(}\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{1-\lambda_{2}}\Big{)}}_{{\mathrm{deterministic\ part}}}$
$+\mathcal{O}\bigg{(}\underbrace{\frac{\alpha\sigma^{2}}{\mu n}+\frac{\chi L\sigma^{2}\alpha^{2}}{\mu(1-\lambda_{2})}}_{{\mathrm{stochastic\ part}}}\bigg{)}.$ (6)
For the non-convex setting, Theorem 1 demonstrates that the ProxSkip algorithm converges to a radius around some stationary point. Without any additional assumptions, a stationary point is the best guarantee possible and is a satisfactory criterion to measure the performance of distributed methods with nonconvex objectives [31]. For the convex case, Theorem 1 shows that ProxSkip converges around some optimal solution. When $\sigma^{2}=0$, i.e., in the deterministic case, ProxSkip converges exactly with sublinear and linear rates for N-CVX/CVX and S-CVX settings, respectively.
Note that stochastic part in convergence rates (1), (1), and (1), which all can be rewritten as $\mathcal{O}(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2})$. It follows from Theorem 1 that, in non-convex, convex, and strongly convex settings, the convergence rate of ProxSkip is
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right),$
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right),$
$\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}\right]\leq\mathcal{O}\left((1-\alpha\mu)^{T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right),$
respectively. Thus, it is established in Theorem 1 that the linear speedup term $\mathcal{O}(\frac{\alpha\sigma^{2}}{n})$ can be achieved. When the stepsize is sufficiently small the term $\mathcal{O}(\frac{\alpha\sigma^{2}}{n})$ dominates convergence rates (1), (1), and (1), which improve linearly with the number of nodes $n$.
Setting $\alpha=\sqrt{\frac{n}{T}}$ for sufficiently large $T$ for non-convex and convex settings, it holds that the rates are bounded by
$\mathcal{O}\left(\sqrt{\frac{1+\sigma^{2}}{nT}}+\frac{n\chi\sigma^{2}}{(1-\lambda_{2})T}+\frac{n\chi^{2}}{(1-\lambda_{2})^{2}T^{2}}\right)=\mathcal{O}\left(\sqrt{\frac{1}{nT}}\right).$
For the strongly convex setting, letting $\alpha=\frac{4\ln T^{2}}{\mu T}$ for sufficiently large $T$, it holds that $1-\frac{\alpha\mu}{4}\leq{\mathrm{exp}}(-\frac{\alpha\mu T}{4})=\frac{1}{T^{2}}$, where ${\mathrm{exp}}(\cdot)$ denotes the exponential function, thus the rate is bounded by
$\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nT}+\frac{1}{T^{2}}+\frac{\chi\sigma^{2}}{(1-\lambda_{2})T^{2}}+\frac{\chi}{(1-\lambda_{2})T^{4}}\right)=\tilde{\mathcal{O}}\left(\frac{1}{nT}\right).$
When $T$ is sufficiently large, the term $\frac{1}{\sqrt{nT}}$ (or $\frac{1}{nT}$ for the strongly convex setting) will dominate the rate. In this scenario, ProxSkip requires $T=\Omega\left(\frac{1}{n\epsilon^{2}}\right)$ (or $T=\Omega\left(\frac{1}{n\epsilon}\right)$) iterations to reach a desired $\epsilon$-accurate solution, thus the convergence accuracy improves linearly with $n$.
In addition, based on Theorem 1, we can even get a tighter rate by carefully selecting the stepsize to obtain the following result.
Corollary 1.
Same settings as in Theorem 1, we have the following convergence results.
Non-convex: It holds that $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\epsilon$ after
$\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p\sqrt{\chi}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{(1-\lambda_{2})\epsilon}\right)$ (7)
expected communication rounds.
Convex: Under the additional Assumption 3 with $\mu\geq 0$, it holds that $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(\bar{{\bf{x}}}^{t})-f^{\star}]\leq\epsilon$ after
$\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p\sqrt{\chi}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{(1-\lambda_{2})\epsilon}\right)$ (8)
expected communication rounds.
Strongly Convex: Under the additional Assumption 3 with $\mu>0$, it holds that $\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]\leq\epsilon$ after
$\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p\sqrt{\chi}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\chi\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{1-\lambda_{2}}\right)$ (9)
expected communication rounds. Here, the notation $\tilde{\mathcal{O}}(\cdot)$ ignores logarithmic factors.
We provide Table II to compare the convergence results of ProxSkip with existing state-of-the-art distributed optimization algorithms, such as local-DSGD [31], $K$-GT [37], and LED [42], with local updates in terms of the number of communication rounds needed to achieve $\epsilon>0$.
TABLE II: A comparison with existing methods employing local steps. $\rho=1-\lambda_{2}$, $K$ denotes the number of local steps, and SL-NI denotes linear speedup with network-independent stepsizes.
Method # communication rounds LS-NI
N-CVX/CVX S-CVX S-CVX
local-DSGD [31] $\mathcal{O}\left(\frac{\sigma^{2}}{nK\epsilon^{2}}+\left(\frac{\sigma}{\sqrt{\rho K}}+\frac{\varsigma}{\rho}\right)\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{\rho\epsilon}\right)$a $\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nK\epsilon}+\left(\frac{\sigma}{\sqrt{\rho K}}+\frac{\varsigma}{\rho}\right)\frac{1}{\sqrt{\epsilon}}+\frac{1}{\rho}\mathrm{log}\nicefrac{{1}}{{\epsilon}}\right)$
[…]
LED [42] $\mathcal{O}\left(\frac{\sigma^{2}}{nK\epsilon^{2}}+\left(\frac{\sigma}{\sqrt{\rho K}}\right)\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{\rho\epsilon}\right)$ $\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nK\epsilon}+\left(\frac{\sigma}{\sqrt{\rho K}}\right)\frac{1}{\sqrt{\epsilon}}+\frac{1}{\rho}\mathrm{log}\nicefrac{{1}}{{\epsilon}}\right)$
ProxSkip $\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p\sqrt{\chi}}{\sqrt{\rho}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{\rho\epsilon}\right)$ $\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p\sqrt{\chi}}{\sqrt{\rho}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\chi}{\rho}\mathrm{log}\nicefrac{{1}}{{\epsilon}}\right)$
  • a
    $\varsigma^{2}$ is function heterogeneity constant such that $\nicefrac{{1}}{{n}}\sum_{i=1}^{n}\|\nabla f_{i}({\bf{x}})-\nabla f({\bf{x}}^{\star})\|^{2}\leq\varsigma^{2}$
  • b
    The results is for the non-convex setting, and no corresponding result is given for the convex setting
Achieving acceleration by $p$ and $n$. According to (7), (8), and (9), when $\epsilon$ is sufficiently small, the convergence rate is dominated by noise and is unaffected by the graph parameter $1-\lambda_{2}$ for ProxSkip. After enough transient time, ProxSkip with $\mathcal{O}\big{(}\frac{p\sigma^{2}}{n\epsilon^{2}}\big{)}$ (or $\tilde{\mathcal{O}}\big{(}\frac{p\sigma^{2}}{n\epsilon}\big{)}$ for the strongly convex setting) achieves linear speedup by the probability of communication $p$ and the number of nodes $n$.
Removing dependence on data heterogeneity. According to Table II, the second term of the communication complexity of local-DSGD [31], a popular algorithm for distributed optimization, is as follows:
N-CVX/CVX: $\mathcal{O}\left(\left(\nicefrac{{\sigma}}{{\sqrt{\rho K}}}+\nicefrac{{\varsigma}}{{\rho}}\right)\epsilon^{-\nicefrac{{3}}{{2}}}\right),$
S-CVX: $\tilde{\mathcal{O}}\left(\left(\nicefrac{{\sigma}}{{\sqrt{\rho K}}}+\nicefrac{{\varsigma}}{{\rho}}\right)\epsilon^{-\nicefrac{{1}}{{2}}}\right).$
Here, $\varsigma^{2}$ represents the function heterogeneity constant such that $\nicefrac{{1}}{{n}}\sum_{i=1}^{n}\|\nabla f_{i}(\mathbf{x})-\nabla f(\mathbf{x}^{\star})\|^{2}\leq\varsigma^{2}$. We note that ProxSkip lacks the additional term $\frac{\varsigma}{\rho}\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}$ ($\frac{\varsigma}{\rho}\frac{1}{\sqrt{\epsilon}}$ for the strongly convex case). Thus, ProxSkip effectively eliminates dependence on the data heterogeneity level $\varsigma^{2}$.
Comparable with existing distributed algorithms incorporating local updates. When $p<\lambda_{2}$, we have $\chi=\mathcal{O}(\max\{1,\frac{1-p}{1-\lambda_{2}}\})=\mathcal{O}(\frac{1}{1-\lambda_{2}})$; when $p>\lambda_{2}$, $\chi=\mathcal{O}(\max\{1,\frac{1-p}{1-\lambda_{2}}\})=\mathcal{O}(1)$. Letting $p=\nicefrac{{1}}{{K}}$ and highlighting the network quantities, the second + third term of the communication complexity of ProxSkip $\mathcal{O}\left(K^{-1}\rho^{-1}+\rho^{-2}\right)$ when $p<1-\rho$, $\mathcal{O}\left(K^{-1}\rho^{-\nicefrac{{1}}{{2}}}+\rho^{-1}\right)$ when $p\geq 1-\rho$. Compared with GT based methods [35, 36, 37], the network dependent bounds are improved. Since the first item of communication complexity for these algorithms is the same, the convergence rates of ProxSkip are comparable with these existing distributed algorithms incorporating local updates.
III-C Linear speedup with network-independent stepsizes
Theorem 2.
Suppose that Assumptions 1, 2, and 4 hold, and $f_{i}$ is $\mu$-strongly convex for some $0<\mu\leq L$. If $0<\alpha\leq\frac{1}{2L}$, $\beta=p$, there exists $\chi=\mathcal{O}(\max\{\frac{1}{p},\frac{1}{1-\lambda_{2}},\frac{1-p}{1-\lambda_{2}}\})$ such that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq\tilde{\zeta}_{0}^{t+1}a_{0}+\frac{\alpha\sigma^{2}}{n\mu}+\mathcal{O}(\alpha^{2}\sigma^{2}),$ (10)
where $a_{0}$ is a constant that depends on the initialization and $\tilde{\zeta}_{0}=\max\left\{1-\alpha\mu,\sqrt{1-\frac{(1-\lambda_{2})p^{2}}{2\chi}}\right\}<1$.
According to this rate, a linear speedup term of $\mathcal{O}(\frac{\alpha\sigma^{2}}{n})+\mathcal{O}(\alpha^{2})$ can be achieved. Importantly, the upper bound on the step size is independent of network topologies, making it a favorable property for practical implementation. Referring to Table II, in the strongly convex setting, while local-DSGD [31], $K$-GT [37], and LED [42] achieve linear speedup bounds, this property hinges on the requirement of network-dependent step sizes, wherein these step sizes are correlated with $1-\lambda_{2}$. In contrast, the step size condition for ProxSkip is $0<\alpha\leq\frac{1}{2L}$, which remains independent of $1-\lambda_{2}$.
Notably, [51] for the first time prove that NIDS/ED/D2 can achieve linear speedup with network-independent stepsizes. However, it remains an open question whether, with local updates, a linear speedup bound can be achieved using network-independent stepsizes. Theorem 2 offers a positive response to this question.
III-D Proof sketch of the main theorem
The existing convergence analysis of ProxSkip [43, 44, 45, 47, 48, 49] relies on primal-dual methodologies. Nevertheless, these analyses are limited to the use of first-order (stochastic) gradient information, leading to a suboptimal exploitation of the available function data. We propose a new proof that, in order to fully utilize function and gradient information, we use matrix factorization techniques to equivalently transform the iteration of ProxSkip into “SGD + consensus” form. Here, we provide a proof sketch for Theorem 1 concerning non-convex objectives. The complete proof of Theorem 1 and Corollary 1 can be found in the Appendix. Here, we provide a proof sketch for Theorem 1 concerning non-convex objectives.
Step 1. (Lemma 1) We first give the equivalent form of update (2) as follows.
$\underbrace{\bar{{\bf{x}}}^{t+1}=\bar{{\bf{x}}}^{t}-\alpha\bar{{\bf{g}}}^{t}}_{\mathrm{SGD}},\quad\underbrace{\mathcal{E}^{t+1}={\bf{\Gamma}}\mathcal{E}^{t}+\bm{\Theta}_{1}^{t}+\bm{\Theta}_{2}^{t}}_{\mathrm{consensus}},$ (11)
where $\bar{{\bf{g}}}^{t}=\frac{1}{n}\sum_{i=1}^{n}{\bf{g}}_{i}^{t}$, $\|{\bf{\Gamma}}\|<1$, $\mathcal{E}^{t}$ measures “consensus”, $\bm{\Theta}_{1}^{t}$ is related to the stochastic gradient, and $\bm{\Theta}_{2}^{t}$ measures “communication error” ($\bm{\Theta}_{2}^{t}=0$, if $\theta_{t}=1$). This description may be less rigorous, but it helps to understand the proof more clearly. See Lemma 1 in Appendix for more details.
Step 2. (Lemma 2) Based on this equivalent update of ProxSkip and by the $L$-smoothness of $f_{i}$, we establish the following descent inequality.
$\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t+1})\right]\leq$ $f(\bar{{\bf{x}}}^{t})-\frac{\alpha}{2}\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}$
$+\frac{2\alpha L^{2}}{n}\|\mathcal{E}^{t}\|^{2}+{\frac{L\alpha^{2}\sigma^{2}}{2n}}.$
Taking average for both sides over $t=0,1,\ldots,T-1$, we have that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]$
$\leq\frac{2F_{0}}{\alpha T}+\frac{4L^{2}}{nT}\sum_{t=0}^{T-1}\|\mathcal{E}^{t}\|^{2}+{\frac{\alpha L\sigma^{2}}{n}}.$ (12)
Step 3. (Lemma 2) We bound $\mathbb{E}\!\left[\|{\bf{\Gamma}}\mathcal{E}^{t}+\bm{\Theta}_{1}^{t}\|^{2}\right]$ and $\mathbb{E}\!\left[\|\bm{\Theta}_{2}^{t}\|^{2}\right]$ by $\|\mathcal{E}^{t}\|^{2}$, $\sigma^{2}$, and $\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}$.
Step 4. (Lemma 2) Subsequently, we establish the following consensus inequality.
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|^{2}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|^{2}+A_{1}n\alpha^{4}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}$
$+A_{2}n\alpha^{2}\sigma^{2}+A_{3}\alpha^{4}\sigma^{2},$
where $10<\tilde{\gamma}<$ 1 and $A_{1},A_{2},A_{3}>0$. Unrolling this recurrence, we have
$\mathbb{E}\!\left[\|\mathcal{E}^{t}\|^{2}\right]\leq$ $\frac{\|\mathcal{E}^{0}\|^{2}}{1-\tilde{\gamma}}+\frac{A_{2}n\alpha^{2}\sigma^{2}+A_{3}\alpha^{4}\sigma^{2}}{1-\tilde{\gamma}}$
$+A_{1}n\alpha^{4}\sum_{k=0}^{t-1}\tilde{\gamma}^{t-1-k}\|\nabla f(\bar{{\bf{x}}}^{k})\|^{2}.$
Step 5. Since
$\sum_{t=0}^{T-1}\sum_{k=0}^{t-1}\tilde{\gamma}^{t-1-k}\|\nabla f(\bar{{\bf{x}}}^{k})\|^{2}\leq\frac{1}{1-\tilde{\gamma}}\sum_{t=0}^{T-1}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2},$
letting $\alpha$ such that
$\frac{4L^{2}A_{1}\alpha^{4}}{1-\tilde{\gamma}}\leq\frac{1}{2},$
it gives that
$\frac{4L^{2}}{nT}\sum_{t=0}^{T-1}\big{\|}\mathcal{E}^{t}\big{\|}_{\mathrm{F}}^{2}\leq\mathcal{O}(\alpha^{2}\sigma^{2})+\frac{1}{2T}\sum_{t=0}^{T-1}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}.$
Combining it with (III-D), we complete the proof, i.e.,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right).$
IV Experiments
We empirically verify the theoretical results of ProxSkip for distributed stochastic optimization. For all experiments, we first compute the solution ${\bf{x}}^{\star}$ or $\|\nabla f({\bf{x}}^{\star})\|^{2}$ to (1) by centralized methods, and then run over a randomly generated connected network with $n$ agents and $\frac{\iota n(n-1)}{2}$ undirected edges, where $\iota$ is the connectivity ratio. The mixing matrix ${\bf{W}}$ is generated with the Metropolis-Hastings rule. All stochastic results are averaged over 10 runs.
IV-A Synthetic Dataset
We consider the linear regression problem on synthetic dataset to show the algorithm performance with data heterogeneity. We construct the distributed least squares objective with $f_{i}({\bf{x}})=\nicefrac{{1}}{{2}}\|{\bf{A}}_{i}{\bf{x}}-{\bf{b}}\|^{2}$ with fixed Hessian ${\bf{A}}_{i}^{2}=\nicefrac{{i^{2}}}{{n^{2}}}{\bf{I}}_{d}$, and sample each ${\bf{b}}_{i}\thicksim\mathcal{N}(0,\nicefrac{{\varsigma^{2}}}{{i^{2}}}{\bf{I}}_{d})$ for each node $i\in[n]$, where $\varsigma^{2}$ can control the deviation between local objectives [31]. Stochastic noise is controlled by adding Gaussian noise with $\sigma^{2}=1$. We use a ring topology with $10$ nodes for this experiment. For all algorithms, we use the same stepsize $\alpha=0.001$.
The results are shown in Fig. 1. According to Fig. 1 the “client-drift” only happens for local-DSGD [31, 22] where the larger $\varsigma^{2}\neq 0$ gets, the poorer model quality D-SGD ends up with. Additionally, the “client-drift” for local-DSGD is even more severe with increasing the number of local updates $\nicefrac{{1}}{{p}}$. However, $K$-GT [37], LED [42], and ProxSkip do not suffer from “client-drift” and ultimately reach the consistent level of model quality regardless of increasing of $\varsigma^{2}$ and $\nicefrac{{1}}{{p}}$. Moreover, from $\nicefrac{{1}}{{p}}=1$ to $\nicefrac{{1}}{{p}}=10$ in Fig. 1, $K$-GT [37], LED [42], and ProxSkip reach the same target after $10000$ rounds to only $1000$ achieving linear speedup in communication by local steps. However, more local steps makes local-DSGD suffer even more in model quality. This is consistent with the theoretical results from [37] and [42].
Figure 1: Learning synthetic convex function over $10$ nodes with noise $\sigma^{2}=1$ (Local-DSGD [22, 31], $K$-GT [37], and LED [42]). All uses the same stepsize and are averaged by ten repetitions. The probability of communication for ProxSkip is $p$, and the number of local updates of local-DSGD, $K$-GT, and LED are $\nicefrac{{1}}{{p}}$.
Figure 2: Experimental results for ProxSkip to logistic regression problem with a strongly convex regularizer $r({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|^{2}$ over ijcnn1 dataset.
Figure 3: Experimental results of logistic regression problem on the ijcnn1 dataset with regularizer $r({\bf{x}})=\frac{L}{200}\|{\bf{x}}\|^{2}$, where $1-\lambda_{2}\approx 0.25$. We set $p=\nicefrac{{1}}{{\sqrt{(1-\lambda_{2})\kappa}}}\approx 0.2$ (the theory predicted optimal choice).
IV-B Real-world Dataset ijcnn1
In this subsection, similar as [43], we also demonstrate our findings on the logistic regression problem with a regularizer. The objective function is
$f({\bf{x}})=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\ln(1+e^{-(\mathcal{A}_{ij}^{\sf T}{{\bf{x}}})\mathcal{B}_{ij}})+r({\bf{x}}).$
Here, $r({\bf{x}})$ is the regularizer, any node $i$ holds its own training date $\left(\mathcal{A}_{ij},\mathcal{B}_{ij}\right)\in$ $\mathbb{R}^{d}\times\{-1,1\},j=1,\cdots,m_{i}$, including sample vectors $\mathcal{A}_{ij}$ and corresponding classes $\mathcal{B}_{ij}$. We use the dataset ijcnn1 from the widely-used LIBSVM library [52], whose attributes is $d=22$ and $\sum_{i=1}^{n}m_{i}=49950$. Moreover, the training samples are randomly and evenly distributed over all the $n$ agents. We control the stochastic noise $\sigma^{2}$ by adding Gaussian noise to every stochastic gradient, i.e., the stochastic gradients are generated as follows: $\nabla F_{i}({\bf{x}})=\nabla f_{i}({\bf{x}})+\omega_{i}$, where $\omega_{i}\thicksim\mathcal{N}(0,\sigma^{2}{\bf{I}}_{d})$ and $\sigma^{2}=10^{-3}$.
Figure 4: Linear speedup of ProxSkip in non-convex settings over ijcnn1 dataset.
Figure 5: Experimental comparison with the same stepsize in non-convex settings over ijcnn1 dataset.
Figure 6: Experimental comparison with tuning stepsize in non-convex settings over ijcnn1 dataset.
1) Strongly Convex Regularizer: We choose the regularizer $r({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|^{2}$ to demonstrate the results in the strongly convex setting. The results are shown in Fig. 2. The relative error $\nicefrac{{\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|}}{{\|{\bf{x}}^{\star}\|}}$ is shown on the $y$-axis. Here, we set $\alpha=\frac{1}{2L}$, which independent of the network topology. We show the performance of ProxSkip at different number of nodes $n$, network connectivity $\iota$, and communication probability $p$. The results show that, when the number of nodes is increased, the relative errors of ProxSkip is reduced under a constant and network-independent stepsize, which validates our results about linear speedup. Moreover, Fig. 2 shows that we can save on communication rounds by reducing $p$, i.e., increasing the number of local steps reduces the amount of communication required to achieve the same level of accuracy. In addition, we compare performance of ProxSkip with Local-DSGD [22, 31], local-GT [35, 36], $K$-GT [37], and LED [42] under stochastic and deterministic settings, respectively, and the results are shown in Fig. 3. In the strongly convex setting, ProxSkip allows for a more relaxed stepsize, i.e., $\alpha=O\left(\frac{1}{L}\right)$, which is independent of the network topology as well as the number of local steps. Instead, the stepsize of the LED is $\alpha=O\left(\frac{\mu(1-\lambda_{2})}{KL^{2}}\right)$. Moreover, when $\sigma^{2}=0$, from Theorem 1, the communication complexity of ProxSkip is $\tilde{O}\left(\sqrt{\frac{\kappa}{1-\lambda_{2}}}\right)$, but LED is $\tilde{O}\left(\frac{\kappa^{2}}{1-\lambda_{2}}\right)$. Thus, ProxSkip would theoretically perform better for the strongly convex setting.
1) Non-convex Regularizer: We choose the regularizer $r({\bf{x}})=\sum_{j=1}^{d}\frac{{\bf{x}}(j)^{2}}{1+{\bf{x}}(j)^{2}}$, to demonstrate the results in the non-convex setting, where ${\bf{x}}=\mathrm{col}\{{\bf{x}}(j)\}_{j=1}^{d}\in\mathbb{R}^{d}$. In this case, we first show that the ProxSkip algorithm achieves linear speedup even in the non-convex setting. In Fig. 4, we conduct experiments at the ring and at $\iota=0.5$, respectively, and the experimental results coincide with the theoretical results of Theorem 1. Setting $n=10$ and $\iota=0.1$, we also compare ProxSkip to the distributed methods Local-DSGD [22, 31], local-GT [35, 36], $K$-GT [37], and LED [42] for different local steps $\nicefrac{{1}}{{p}}=10,5,1$. We use the same stepsize $\alpha=0.01$ for all algorithms. From Fig. 5 and 6, it shows that ProxSkip and LED perform similarly, and they outperforms the other methods as we increase the number of local steps.
V Conclusion
This paper revisits the convergence bounds of ProxSkip for distributed stochastic optimization. We present a new analysis with a novel proof technique applicable to stochastic non-convex, convex, and strongly convex settings. Through this comprehensive analysis, we derive several new results that rival the bounds of state-of-the-art distributed algorithms [31, 35, 36, 37, 42]. We establish that the leading communication complexity of ProxSkip is $\mathcal{O}(pn^{-1}\sigma^{2})$, indicating that ProxSkip can achieve acceleration by $p$ and $n$. Our proposed proof technique overcomes the analytical limitations of prior work [43, 44, 45, 46, 47, 48, 49] and might be of independent interest in the community.
Appendix
I Preliminaries
I-A Basic Facts
The stochastic processes such as randomized communication and gradient estimation generate two sequences of $\sigma$-algebra. We denote by $\mathcal{G}^{t}$ the $\sigma$-algebra of gradient estimation at $t$-th iteration and $\mathcal{F}^{t}$ the $\sigma$-algebra of randomized communication at the same step. The sequences $\{\mathcal{G}^{t}\}_{t\geq 0}$ and $\{\mathcal{F}^{t}\}_{t\geq 0}$ satisfy
$\mathcal{G}^{0}\subset\mathcal{F}^{0}\subset\mathcal{G}^{1}\subset\mathcal{F}^{1}\subset\mathcal{G}^{2}\subset\mathcal{F}^{2}\subset\cdots\subset\mathcal{G}^{t}\subset\mathcal{F}^{t}\subset\cdots\ .$
With these notations, we can clarify the stochastic dependencies among the variables generated by Algorithmd 1, i.e., $({\bf{G}}^{t},\hat{{\bf{Z}}}^{t})$ is measurable in $\mathcal{G}^{t}$ and $({\bf{Y}}^{t+1},{\bf{X}}^{t+1})$ is measurable in $\mathcal{F}^{t}$.
The Bregman divergence of $f$ at points $(x,y)$ is defined by
$D_{f}(x,y):=f(x)-f(y)-\langle\nabla f(y),x-y\rangle.$
It is easy to verify that $\langle\nabla f(x)-\nabla f(y),x-y\rangle=D_{f}(x,y)+D_{f}(y,x)$. If $f$ is convex, from the definition of convex function, we have $D_{f}(x,y)\geq 0$ and $D_{f}(y,x)\geq 0$. Thus
$\langle\nabla f(x)-\nabla f(y),x-y\rangle\geq D_{f}(x,y),\text{ and }\langle\nabla f(x)-\nabla f(y),x-y\rangle\geq D_{f}(y,x).$ (13)
For an $L$-smooth and $\mu$-strongly convex function $f$, by [43, Appendix. A] we have
$\frac{\mu}{2}\|x-y\|^{2}\leq$ $D_{f}(x,y)\leq\frac{L}{2}\|x-y\|^{2},$ (14)
$\frac{1}{2L}\|\nabla f(x)-\nabla f(y)\|^{2}\leq$ $D_{f}(x,y)\leq\frac{1}{2\mu}\|\nabla f(x)-\nabla f(y)\|^{2}.$ (15)
Under the $L$-smoothness condition, we have
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|x-y\|^{2},\ \forall x,y\in\mathbb{R}^{d}$ (16)
I-B Notations
For any $n\times m$ matrices $\mathbf{a}$ and $\mathbf{b}$, their inner product is denoted as $\langle\mathbf{a},\mathbf{b}\rangle=\mathrm{Trace}(\mathbf{a}^{\sf T}\mathbf{b})$. For a given matrix $\mathbf{a}$, the Frobenius norm is given by $\|\mathbf{a}\|_{\mathrm{F}}$, while the spectral norm is given by $\|\mathbf{a}\|$. Define the gradient and communication noise as
gradient noise: ${\bf{S}}^{t}=[{\bf{s}}^{t}_{1},\ldots,{\bf{s}}^{t}_{n}]^{\sf T}={\bf{G}}^{t}-\nabla F({\bf{X}}^{t}),\text{ where }{\bf{s}}^{t}_{i}={\bf{g}}_{i}^{t}-\nabla f_{i}({\bf{x}}_{i}^{t});$
communication noise: ${\bf{E}}^{t}=\frac{\theta_{t}-1}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}.$
We also define the following notations to simplify the analysis:
$\bar{{\bf{x}}}^{t}\triangleq\big{(}\frac{1}{n}\sum_{i=1}^{n}{\bf{x}}_{i}^{t}\big{)}^{\sf T},\quad\bar{{\bf{X}}}^{t}={\bf{1}}\otimes\bar{{\bf{x}}}^{t},\quad\bar{{\bf{s}}}^{t}\triangleq\big{(}\frac{1}{n}\sum_{i=1}^{n}{\bf{s}}_{i}^{t}\big{)}^{\sf T},\quad\overline{\nabla F}({\bf{X}}^{t})\triangleq\big{(}\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}({\bf{x}}^{t}_{i})\big{)}^{\sf T}.$
With Assumption 1 and [5, Section IV-B], the mixing matrix ${\bf{W}}$ can be decomposed as
${\bf{W}}={\bf{P\Lambda P}}^{-1}=\left[\begin{array}[]{cc}{\bf{1}}&\hat{{\bf{P}}}\\ \end{array}\right]\left[\begin{array}[]{cc}{\bf{I}}&0\\ 0&\hat{{\bf{\Lambda}}}\\ \end{array}\right]\left[\begin{array}[]{c}\frac{1}{n}{\bf{1}}^{\sf T}\\ \hat{{\bf{P}}}^{\sf T}\\ \end{array}\right],$
where $\hat{{\bf{\Lambda}}}=\mathrm{diag}\{\lambda_{2},\ldots,\lambda_{n}\}$, and matrix $\hat{{\bf{P}}}\in\mathbb{R}^{n\times(n-1)}$ satisfies
[…]
Therefore, the matrix ${\bf{W}}_{a}$ and ${\bf{W}}_{b}$ can be decomposed as
${\bf{W}}_{a}=\left[\begin{array}[]{cc}{\bf{1}}&\hat{{\bf{P}}}\\ \end{array}\right]\underbrace{\left[\begin{array}[]{cc}{\bf{1}}&0\\ 0&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]}_{:={\bf{\Lambda}}_{a}}\left[\begin{array}[]{c}\frac{1}{n}{\bf{1}}^{\sf T}\\ \hat{{\bf{P}}}^{\sf T}\\ \end{array}\right],~{}{\bf{W}}_{b}^{2}=\left[\begin{array}[]{cc}{\bf{1}}&\hat{{\bf{P}}}\\ \end{array}\right]\underbrace{\left[\begin{array}[]{cc}0&0\\ 0&\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]}_{:={\bf{\Lambda}}_{b}^{2}}\left[\begin{array}[]{c}\frac{1}{n}{\bf{1}}^{\sf T}\\ \hat{{\bf{P}}}^{\sf T}\\ \end{array}\right],$ (27)
where $\hat{{\bf{\Lambda}}}_{a}={\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})$, $\hat{{\bf{\Lambda}}}_{b}=\sqrt{{\bf{I}}-\hat{{\bf{\Lambda}}}}$. Since $\lambda_{i}\in(-1,1)$ for $i=2,\dots,n$, it holds that $1-\frac{1}{2\chi}(1-\lambda_{i})\in[0,1)$ and $0\preceq{\bf{W}}_{a}\prec{\bf{I}}$ for $\chi\geq 1$.
II Proof of Theorem 1 and Corollary 1
II-A Transformation and Some Descent Inequalities
Here, we introduce an auxiliary variable ${\bf{R}}^{t}={\bf{Y}}^{t}+\alpha\nabla F(\bar{{\bf{X}}}^{t})$, where $\bar{{\bf{X}}}^{t}=\mathbf{1}\otimes\bar{{\bf{x}}}^{t}$. It follows from (2b) and (2c) that, when $\beta=1$ and $p=1$, ${\bf{Y}}^{t+1}={\bf{Y}}^{t}+\frac{1}{2\chi}{\bf{W}}_{b}^{2}\hat{{\bf{Z}}}^{t}$. For any fixed point $({\bf{X}},{\bf{Y}})$ of update (2), it holds that $\hat{{\bf{Z}}}={\bf{X}}$, ${\bf{Y}}+\alpha\nabla F({{\bf{X}}})=0$, ${\bf{W}}_{b}{\bf{X}}=0$. Thus, ${\bf{R}}=0$ implies that $\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}({\bf{x}})=0$, i.e., ${\bf{x}}$ is a stationary point of problem (1). By this new variable, we give following error dynamics of Algorithm 1.
Lemma 2.
Suppose that Assumption 1 holds. There exists a invertible matrix ${\bf{Q}}$ and a diagonal matrix ${\bf{\Gamma}}$ such that
$\bar{{\bf{x}}}^{t+1}$ $=\bar{{\bf{x}}}^{t}-\alpha\overline{\nabla F}({\bf{X}}^{t})-\alpha\bar{{\bf{s}}}^{t},$ (28a)
$\mathcal{E}^{t+1}$ $=\underbrace{{\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}+\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]}_{:=\mathbb{G}^{t}}+\underbrace{{\bf{Q}}^{-1}\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right]}_{:=\mathbb{F}^{t}},$ (28f)
where $\gamma\triangleq\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1$,
$\mathcal{E}^{t}\triangleq{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{{\bf{R}}}^{t}\\ \end{array}\right],\quad\left\{\begin{array}[]{l}{\bf{\Sigma}}_{1}^{t}=\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t},\\ {\bf{\Sigma}}_{2}^{t}=\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1})\end{array}\right..$
Moreover, we have
$\|{\bf{Q}}\|^{2}\leq 2\text{ and }\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}.$
In addition, we have
$\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}.$ (29)
Proof See Appendix IV. $\Box$
Based on Lemma 2, we give the following descent inequalities.
Lemma 3.
Suppose that Assumptions 1, 2, and 4 hold. If $\alpha\leq\frac{1}{2L}$, it holds that
$\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t+1})\;|\;\mathcal{G}^{t}\right]\leq$ $f(\bar{{\bf{x}}}^{t})-\frac{\alpha}{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha L^{2}}{n}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n},$ (30)
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}},$ (31)
where
$\tilde{\gamma}=\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2(1-p)\big{(}3+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\big{)}}{\chi^{2}}.$ (32)
Moreover, if $f_{i}$ is $\mu$-convex ($\mu\geq 0$) and $\alpha\leq\frac{1}{4L}$, it holds that
$\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\|^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq(1-\mu\alpha)\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-\alpha(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})),$ (33)
where ${\bf{x}}^{\star}$ solves problem (1).
Proof See Appendix V. $\Box$
II-B Convergence Analysis: Non-convex
With Lemma 2 and Lemma 3, we further have the following theorem.
Theorem 3.
Suppose that Assumptions 1, 2, and 4 hold. If $\beta=1$, $\alpha$ and $\chi$ satisfy that $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},$ (34)
it holds that $\tilde{\gamma}\leq\frac{1+\gamma}{2}<1$ and
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]\leq$ $\frac{4(f(\bar{{\bf{x}}}^{0})-f^{\star})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}\big{(}\chi^{3}+256\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}},$ (35)
where $\varsigma^{2}_{0}=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\bar{{\bf{x}}}^{0})-\nabla f(\bar{{\bf{x}}}^{0})\|^{2}$.
Proof See Appendix VI. $\Box$
Based on Theorem 3, we can even get a tighter rate by carefully selecting the stepsize similar to [31, Lemma 17], [37, Lemma C.13], or [42, Corollary 1].
Corollary 2.
Suppose that Assumptions 1, 2, and 4 hold. If $\beta=1$, $\chi$ satisfies that $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right)$ such that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{(1-\lambda_{2})T}\right).$ (36)
Proof See Appendix VII. $\Box$ When $p<\lambda_{2}$, we have $\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}=\mathcal{O}\left(\frac{1}{1-\lambda_{2}}\right)$. Choosing $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$. Since in each iteration we trigger communication with probability $p$, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\epsilon$ is bounded by:
$p<\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{1-\lambda_{2}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{(1-\lambda_{2})^{2}}\frac{1}{\epsilon}\right).$
When $p\geq\lambda_{2}$, we have $\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}=\mathcal{O}(1)$. If we choose $\chi$ such that $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, then for any desired accuracy $\epsilon>0$, the expected communication complexity is bounded by
$p\geq\lambda_{2}:\ p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{1-\lambda_{2}}\frac{1}{\epsilon}\right).$
II-C Convergence Analysis: Convex
By Lemma 2 and Lemma 3, we also can deduce the following lemma.
Theorem 4.
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu\geq 0$, if $\beta=1$, $\alpha$ and $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\right\},$ (37)
it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{L\alpha^{2}\sigma^{2}\big{(}\chi^{3}+384\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}}.$ (38)
Proof See Appendix VIII. $\Box$
Similar as the analysis of non-convex setting, with Theorem 4, we have the following results.
Corollary 3.
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu\geq 0$, if $\beta=1$ and $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right)$ such that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{(1-\lambda_{2})T}\right).$ (39)
Proof See Appendix IX. $\Box$ Similar as the analysis of non-convex setting, when $p<\lambda_{2}$ and choosing $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}\leq\mathcal{O}\left(\frac{1}{1-\lambda_{2}}\right)$, for any desired accuracy $\epsilon>0$, the expected communication complexity is bounded by
$p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{1-\lambda_{2}}\frac{\sigma\sqrt{L}}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{(1-\lambda_{2})^{2}}\frac{L}{\epsilon}\right).$
When $p\geq\lambda_{2}$, we have $\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}=\mathcal{O}(1)$. If we choose $\chi$ such that $\chi=\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, then for any desired accuracy $\epsilon>0$, the expected communication complexity is bounded by
$p\times\text{(iteration complexity)}=\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p}{\sqrt{1-\lambda_{2}}}\frac{\sigma\sqrt{L}}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{1-\lambda_{2}}\frac{L}{\epsilon}\right).$
II-D Convergence Analysis: Strongly Convex
By Lemma 2 and Lemma 3, we also can deduce the following lemma.
Theorem 5.
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu>0$, If $\beta=1$, $\alpha$ and $\chi$ satisfy that $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$ and
$0<\alpha\leq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\frac{72\mu}{L^{2}},\frac{1-\gamma}{12L+\nicefrac{{\mu}}{{2}}},\frac{\sqrt[3]{4\mu(1-\gamma)}}{L}\right\},$ (40)
where $\gamma=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1$, it holds that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq$ $\Big{(}1-\frac{\alpha\mu}{4}\Big{)}^{t}\Big{(}\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{1-\lambda_{2}}\Big{)}+\frac{2\alpha\sigma^{2}}{\mu n}+\frac{7L\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{12\mu(1-\lambda_{2})\chi}.$ (41)
Proof See Appendix X. $\Box$
Based on Theorem 5, we can even get a tighter rate by carefully selecting the step size similar to [31].
Corollary 4.
Suppose that Assumptions 1, 2, and 4 hold. Under the additional Assumption 3 with $\mu>0$, if $\beta=1$ and $\chi\geq\max\left\{\frac{288(1-p)}{1-\lambda_{2}},1\right\}$, there exist a constant $\alpha=\mathcal{O}\left(\frac{\mu(1-\lambda_{2})}{\chi L^{2}}\right)$ such that
$\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]\leq\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nT}+\frac{\sigma^{2}\chi}{(1-\lambda_{2})T^{2}}+\mathrm{exp}\Big{[}-\frac{(1-\lambda_{2})T}{\chi}\Big{]}\right).$ (42)
Proof See Appendix XI. $\Box$
Similar as the analysis of non-convex and convex settings, we have $\chi=\max\{\nicefrac{{288(1-p)}}{{1-\lambda_{2}}},1\}\leq\mathcal{O}(\nicefrac{{1}}{{(1-\lambda_{2})}})$ if $p<\lambda_{2}$ and $\chi=\max\{\nicefrac{{288(1-p)}}{{1-\lambda_{2}}},1\}=\mathcal{O}(1)$ if $p\geq\lambda_{2}$. Thus, for any desired accuracy $\epsilon>0$, the expected number of communication rounds required to achieve $\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]\leq\epsilon$ is bounded by
$p\times\text{(iteration complexity)}=\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p}{{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{(1-\lambda_{2})^{2}}\right),\ p\in(0,\lambda_{2}),$
and
$p\times\text{(iteration complexity)}=\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{1-\lambda_{2}}\right),\ p\in[\lambda_{2},1].$
III Proof Theorem 2
Then, we further prove ProxSkip can achieve linear speedup with network-independent stepsize. We introduce new iterates $\{{\bf{U}}^{t}\}$ to facilitate the analysis. Similar techniques can be found, e.g., in [38, 39, 6], ${\bf{Y}}^{t}=\alpha{\bf{W}}_{b}{\bf{U}}^{t}$. Since ${\bf{I}}-{\bf{W}}_{a}=\frac{1}{2\chi}{\bf{W}}^{2}_{b}$, from (2b) and (2c), we have
$\left\{\begin{array}[]{rl}{\bf{X}}^{t+1}\ =&(1-\theta_{t})\hat{{\bf{Z}}}^{t}+\theta_{t}{\bf{W}}_{a}\hat{{\bf{Z}}}^{t}\\ \alpha{\bf{W}}_{b}{\bf{U}}^{t+1}\ =&\alpha{\bf{W}}_{b}{\bf{U}}^{t}+\beta(\hat{{\bf{Z}}}^{t}-{\bf{X}}^{t+1})\end{array}\right.\Longleftrightarrow\left\{\begin{array}[]{rl}{\bf{W}}_{b}{\bf{U}}^{t+1}\ =&{\bf{W}}_{b}{\bf{U}}^{t}+\frac{\beta\theta_{t}}{2\chi\alpha}{\bf{W}}_{b}^{2}\hat{{\bf{Z}}}^{t}\\ {\bf{X}}^{t+1}\ =&\hat{{\bf{Z}}}^{t}-\frac{\alpha}{\beta}{\bf{W}}_{b}({\bf{U}}^{t+1}-{\bf{U}}^{t})\end{array}\right.$
Therefore, letting ${\bf{Y}}^{0}=0$, we have the following equivalent form of ProxSkip (2) in the sense that they generate an identical sequence $({\bf{X}}^{t},\hat{{\bf{Z}}}^{t})$.
$\hat{{\bf{Z}}}^{t}$ $=\ {\bf{X}}^{t}-\alpha{\bf{G}}^{t}-\alpha{\bf{W}}_{b}{\bf{U}}^{t},$ (43a)
${\bf{U}}^{t+1}$ $=\ {\bf{U}}^{t}+\frac{\beta\theta_{t}}{2\chi\alpha}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t},$ (43b)
${\bf{X}}^{t+1}$ $=\ \hat{{\bf{Z}}}^{t}-\frac{\alpha}{\beta}{\bf{W}}_{b}({\bf{U}}^{t+1}-{\bf{U}}^{t}).$ (43c)
This equivalent form is more useful for the subsequent convergence analysis. The optimality condition of problem (1) is as the following lemma.
Lemma 4.
Suppose that Assumption 1 holds. If there exists a point $({\bf{X}}^{\star},{\bf{U}}^{\star})$ that satisfies:
00 0 $=\nabla F({\bf{X}}^{\star})+{\bf{W}}_{b}{\bf{U}}^{\star},$ (44a)
00 0 $={\bf{W}}_{b}{\bf{Z}}^{\star},$ (44b)
then it holds that ${\bf{X}}^{\star}=[{\bf{x}}^{\star},{\bf{x}}^{\star},\ldots,{\bf{x}}^{\star}]^{\sf T}$, where ${\bf{x}}^{\star}\in\mathbb{R}^{d}$ is a stationary point to problem (1).
From Lemma 4, when ${\bf{G}}^{t}=\nabla F({\bf{X}}^{t})$, we have that any fixed point of (43) satisfies the condition (44). We also define the following notations to simplify the analysis:
$\widetilde{{\bf{Z}}}^{t}\triangleq\hat{{\bf{Z}}}^{t}-{\bf{X}}^{\star},\quad\widetilde{{\bf{X}}}^{t}\triangleq{\bf{X}}^{t}-{\bf{X}}^{\star},\quad\widetilde{{\bf{U}}}^{t}\triangleq\alpha({\bf{U}}^{t}-{\bf{U}}^{\star}),\quad\bar{{\bf{e}}}^{t}\triangleq\bar{{\bf{x}}}^{t}-({\bf{x}}^{\star})^{\sf T},$
where $({\bf{X}}^{\star},{\bf{U}}^{\star})$ satisfies (44). Similar as Lemma 2, we give another error dynamics of ProxSkip, which will be used for proving the linear speedup with network-independent stepsizes of ProxSkip under strongly convexity.
Lemma 5.
Suppose that Assumption 1 holds. If $\beta=p$ and $\chi p\geq 1$, there exist a invertible matrix ${\bf{Q}}^{\mathrm{s}}$ and a diagonal matrix ${\bf{\Gamma}}$ such that
$\bar{{\bf{e}}}^{t+1}$ $=\bar{{\bf{e}}}^{t}-\alpha\overline{\nabla F}({\bf{X}}^{t})-\alpha\bar{{\bf{s}}}^{k},$ (45a)
$\mathcal{E}_{\mathrm{s}}^{t+1}$ $=\underbrace{{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\alpha\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]}_{:=\mathbb{G}_{\mathrm{s}}^{t}}+\underbrace{\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ p\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right]}_{:=\mathbb{F}_{\mathrm{s}}^{t}},$ (45f)
where $\upsilon$ is an arbitrary strictly positive constant,
$\mathcal{E}_{\mathrm{s}}^{t}\triangleq\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]\text{ and }\gamma\triangleq\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1.$
Moreover, we have $\|{\bf{Q}}^{\mathrm{s}}\|^{2}\|({\bf{Q}}^{\mathrm{s}})^{-1}\|^{2}\leq\frac{8\chi^{2}}{p^{2}(1+\lambda_{n})}$.
Proof See Appendix XII. $\Box$ With this error dynamics, similar as Lemma 3, we give the following descent inequalities.
Lemma 6.
Suppose that Assumptions 1, 2, and 4 hold, and $f_{i}$ is $\mu$-strongly convex for some $0<\mu\leq L$. Let $\upsilon=1/\|({\bf{Q}}^{\mathrm{s}})^{-1}\|$. If $\alpha\leq\frac{1}{2L}$, it holds that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq(1-\mu\alpha)\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{2\alpha L\vartheta_{\mathrm{s}}}{n}\|\mathcal{E}_{s}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n},$ (46)
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\tilde{\gamma}_{\mathrm{s}}\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+D_{1}\|\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+D_{2}n\alpha^{2}\sigma^{2},$ (47)
where $\vartheta_{\mathrm{s}}=\|{\bf{Q}}^{\mathrm{s}}\|^{2}\|({\bf{Q}}^{\mathrm{s}})^{-1}\|^{2}$, $\tilde{\gamma}_{\mathrm{s}}=\gamma+\frac{3(1-p)(2+p^{2})}{\chi^{2}}$
$D_{1}=\frac{\alpha^{2}L^{2}(2\chi^{2}+p^{2})}{2\chi^{2}(1-\gamma)}+\frac{3\alpha^{2}L^{2}(1-p)(2+p^{2})}{2\chi^{2}},\ D_{2}=\frac{(1-p)(2+p^{2})+(p^{2}+2\chi^{2})}{2\chi^{2}}.$
Proof See Appendix XIII. $\Box$
With Lemmas 5 and 6, we have the following Theorem.
Theorem 6.
Suppose that Assumptions1, 2, and 4 hold, and $f_{i}$ is $\mu$-strongly convex for some $0<\mu\leq L$. If $0<\alpha\leq\frac{1}{2L}$, $\beta=p$, and
$\chi>\max\left\{\frac{1}{p},\frac{36}{1-\lambda_{2}},\frac{72(1-p)}{1-\lambda_{2}}\right\},$ (48)
it holds that $\tilde{\gamma}_{\mathrm{s}}<1$ and
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq\tilde{\zeta}_{0}^{t+1}a_{0}+\mathcal{O}\left(\frac{\alpha^{4}\sigma^{2}L^{3}\chi^{4}}{\mu p^{2}(1-\lambda_{2})^{2}(1-\zeta)}+\frac{\alpha^{2}\sigma^{2}L\chi^{3}}{\mu p^{2}(1-\lambda_{2})}\right)+\frac{\alpha\sigma^{2}}{n\mu},$ (49)
where $a_{0}$ is a constant that depends on the initialization and $\tilde{\zeta}_{0}=\max\{1-\alpha\mu,\sqrt{1-\frac{(1-\lambda_{2})p^{2}}{2\chi}}\}<1$.
Proof See Appendix XIV. $\Box$
IV Proof of Lemma 2
Proof It follows from (2b), $\mathbf{I}-\mathbf{W}_{a}=\frac{1}{2\chi}\mathbf{W}_{b}^{2}$ and ${\bf{E}}^{t}=\frac{1}{2\chi}(\theta_{t}-1){\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$ that
$\mathbf{X}^{t+1}$ $=(1-\theta_{t})\hat{\mathbf{Z}}^{t}+\theta_{t}\mathbf{W}_{a}\hat{\mathbf{Z}}^{t}$
$={\bf{W}}_{a}\hat{\mathbf{Z}}^{t}+(1-\theta_{t})\hat{\mathbf{Z}}^{t}+\theta_{t}\mathbf{W}_{a}\hat{\mathbf{Z}}^{t}-{\bf{W}}_{a}\hat{\mathbf{Z}}^{t}$
$={\bf{W}}_{a}\hat{\mathbf{Z}}^{t}+(1-\theta_{t})({\bf{I}}-{\bf{W}}_{a})\hat{\mathbf{Z}}^{t}$
$={\bf{W}}_{a}\hat{\mathbf{Z}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t}.$
Since $\beta=1$, it follows from (2b), (2c), and $\mathbf{I}-\mathbf{W}_{a}=\frac{1}{2\chi}\mathbf{W}_{b}^{2}$ that
${\bf{Y}}^{t+1}={\bf{Y}}^{t}+\frac{1}{2\chi}{\bf{W}}_{b}^{2}\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}={\bf{Y}}^{t}+({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}.$
By ${\bf{R}}^{t}={\bf{Y}}^{t}+\alpha\nabla F(\bar{{\bf{X}}}^{t})$, ${\bf{\Sigma}}_{2}^{t}=\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1})$, and ${\bf{E}}^{t}=\frac{1}{2\chi}(\theta_{t}-1){\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, we have
${\bf{R}}^{t+1}-{\bf{R}}^{t}$ $={\bf{Y}}^{t+1}-{\bf{Y}}^{t}+\alpha(\nabla F(\bar{{\bf{X}}}^{t+1})-\nabla F(\bar{{\bf{X}}}^{t}))$
$=({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}+\alpha(\nabla F(\bar{{\bf{X}}}^{t+1})-\nabla F(\bar{{\bf{X}}}^{t}))$
$=({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}+{\bf{W}}_{b}{\bf{E}}^{t}-\alpha{\bf{\Sigma}}_{2}^{t}.$
Note that ${\bf{\Sigma}}_{1}^{t}=\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t}$. Algorithm 1 (update (2)) is equivalent to
$\hat{{\bf{Z}}}^{t}$ $={\bf{X}}^{t}-{\bf{R}}^{t}-\alpha{\bf{\Sigma}}_{1}^{t},$
${\bf{X}}^{t+1}$ $={\bf{W}}_{a}\hat{{\bf{Z}}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t},$
${\bf{R}}^{t+1}$ $={\bf{R}}^{t}+({\bf{I}}-{\bf{W}}_{a})\hat{{\bf{Z}}}^{t}-\alpha{\bf{\Sigma}}_{2}^{t}+{\bf{W}}_{b}{\bf{E}}^{t},$
which also can be rewritten as (since ${\bf{W}}_{a}={\bf{I}}-\frac{1}{2\chi}{\bf{W}}^{2}_{b}$)
$\left[\begin{array}[]{c}{\bf{X}}^{t+1}\\ {\bf{R}}^{t+1}\\ \end{array}\right]=$ $\left[\begin{array}[]{cc}{\bf{W}}_{a}&-{\bf{W}}_{a}\\ {\bf{I}}-{\bf{W}}_{a}&{\bf{W}}_{a}\\ \end{array}\right]\left[\begin{array}[]{c}{\bf{X}}^{t}\\ {\bf{R}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}{\bf{W}}_{a}{\bf{\Sigma}}_{1}^{t}\\ \frac{1}{2\chi}{\bf{W}}_{b}^{2}{\bf{\Sigma}}_{1}^{t}+{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]+\left[\begin{array}[]{c}-{\bf{W}}_{b}{\bf{E}}^{t}\\ {\bf{W}}_{b}{\bf{E}}^{t}\\ \end{array}\right].$
Multiplying both sides of the above by $\mathrm{diag}\{{\bf{P}}^{-1},{\bf{P}}^{-1}\}$ on the left, and using (27) and
${\bf{P}}^{-1}{\bf{X}}^{t}=\left[\begin{array}[]{c}\bar{{\bf{x}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t}\\ \end{array}\right],\ {\bf{P}}^{-1}{\bf{R}}^{t}=\left[\begin{array}[]{c}\alpha\overline{\nabla F}(\bar{{\bf{X}}}^{t})\\ \hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t}\\ \end{array}\right],\ {\bf{P}}^{-1}\nabla F({\bf{X}}^{t})=\left[\begin{array}[]{c}\overline{\nabla F}({\bf{X}}^{t})\\ \hat{{\bf{P}}}^{\sf T}\nabla F({\bf{X}}^{t})\\ \end{array}\right],\ {\bf{P}}^{-1}{\bf{E}}^{t}=\left[\begin{array}[]{c}0\\ \hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
we have
$\bar{{\bf{x}}}^{t+1}=$ $~{}\bar{{\bf{x}}}^{t}-\alpha\overline{\nabla F}({\bf{X}}^{t})-\alpha\bar{{\bf{s}}}^{t},$
$\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t+1}\\ \hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t+1}\\ \end{array}\right]=$ $\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\\ {\bf{I}}-\hat{{\bf{\Lambda}}}_{a}&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}+\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
Let
${\bf{H}}=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\\ {\bf{I}}-\hat{{\bf{\Lambda}}}_{a}&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]=\left[\begin{array}[]{cc}{\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})&-({\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}}))\\ \frac{1}{2\chi}({{\bf{I}}-\hat{{\bf{\Lambda}}}})&{\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})\\ \end{array}\right],$
where $\hat{{\bf{\Lambda}}}=\mathrm{diag}\{\lambda_{2},\ldots,\lambda_{n}\}$, and $\lambda_{i}\in(-1,1)$. Since the blocks of ${\bf{H}}$ are diagonal matrices, there exists a permutation matrix ${\bf{Q}}_{1}$ such that ${\bf{Q}}_{1}{\bf{H}}{\bf{Q}}_{1}^{\sf T}=\mathrm{blkdiag}\{H_{i}\}_{i=2}^{n}$, where
$H_{i}=\left[\begin{array}[]{cc}1-\frac{1}{2\chi}(1-\lambda_{i})&-(1-\frac{1}{2\chi}(1-\lambda_{i}))\\ \frac{1}{2\chi}(1-\lambda_{i})&1-\frac{1}{2\chi}(1-\lambda_{i})\\ \end{array}\right].$
Setting $\nu_{i}=1-\frac{1}{2\chi}(1-\lambda_{i})$, we have $\nu_{i}\in(0,1)$ and $H_{i}$ can be rewritten as
$H_{i}=\left[\begin{array}[]{cc}\nu_{i}&-\nu_{i}\\ 1-\nu_{i}&\nu_{i}\\ \end{array}\right]\in\mathbb{R}^{2\times 2}.$
It holds that $\mathrm{Tr}(H_{i})=2\nu_{i}$, $\mathrm{det}(H_{i})=\nu_{i}$. Thus, the eigenvalues of $H_{i}$ are
$\gamma_{(1,2),i}$ $=\frac{1}{2}\Big{[}\mathrm{Tr}(H_{i})\pm\sqrt{\mathrm{Tr}(H_{i})^{2}-4\mathrm{det}(H_{i})}\Big{]}=\nu_{i}\pm\sqrt{\nu_{i}^{2}-\nu_{i}}.$
Notice that $|\gamma_{(1,2),i}|<1$ when $-\nicefrac{{1}}{{3}}<\nu_{i}<1$, which holds under Assumption 1 since ${\bf{W}}_{a}\succ 0$, i.e., $0<\nu_{i}<1\ (i=2,\ldots,n)$. For $10<\nu_{i}<$ 1, the eigenvalues of $H_{i}$ are complex and distinct:
$\gamma_{(1,2),i}=\nu_{i}\pm j\sqrt{\nu_{i}-\nu_{i}^{2}},\ |\gamma_{(1,2),i}|<1,$
where $j^{2}=-1$. Through algebraic multiplication it can be verified that $H_{i}=Q_{2,i}\Gamma_{i}Q_{2,i}^{-1}$, where $\Gamma_{i}=\mathrm{diag}\{\gamma_{1,i},\gamma_{2,i}\}$ and
$Q_{2,i}=\left[\begin{array}[]{cc}\sqrt{\nu_{i}}&\sqrt{\nu_{i}}\\ -j\sqrt{1-\nu_{i}}&j\sqrt{1-\nu_{i}}\end{array}\right],\quad Q_{2,i}^{-1}=\left[\begin{array}[]{cc}\frac{1}{2\sqrt{\nu_{i}}}&\frac{j}{2\sqrt{1-\nu_{i}}}\\ \frac{1}{2\sqrt{\nu_{i}}}&-\frac{j}{2\sqrt{1-\nu_{i}}}\end{array}\right].$
Note that
$Q_{2,i}Q_{2,i}^{*}=\left[\begin{array}[]{cc}2\nu_{i}&0\\ 0&2(1-\nu_{i})\end{array}\right],\text{ and }(Q_{2,i}^{-1})(Q_{2,i}^{-1})^{*}=\frac{1}{4\nu_{i}(1-\nu_{i})}\left[\begin{array}[]{cc}1&1-2\nu_{i}\\ 1-2\nu_{i}&1\end{array}\right].$
Since the spectral radius of matrix is upper bounded by any of its norm and $10<\nu_{i}<$ 1, it holds that
$\|Q_{2,i}\|^{2}\leq\|Q_{2,i}Q_{2,i}^{*}\|_{1}\leq 2,\text{ and }\|Q_{2,i}^{-1}\|^{2}\leq\|(Q_{2,i}^{-1})(Q_{2,i}^{-1})^{*}\|_{1}\leq\frac{2}{4\nu_{i}(1-\nu_{i})}.$
Using $\nu_{i}\geq 1-\frac{1}{2\chi}(1-\lambda_{n})$ and $1-\nu_{i}=\frac{1}{2\chi}(1-\lambda_{i})\geq\frac{1}{2\chi}(1-\lambda_{2})$, we have
$\|Q_{2,i}^{-1}\|^{2}\leq\frac{\chi}{(1-\frac{1}{2\chi}(1-\lambda_{n}))(1-\lambda_{2})}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}\ .$
Let ${\bf{Q}}={\bf{Q}}_{1}^{\sf T}{\bf{Q}}_{2}$ with ${\bf{Q}}_{2}=\mathrm{blkdiag}\{Q_{2,i}\}_{i=2}^{n}$. We have ${\bf{Q}}^{-1}{\bf{H}}{\bf{Q}}={\bf{\Gamma}}$, where ${\bf{\Gamma}}=\mathrm{blkdiag}\{\Gamma_{i}\}_{i=2}^{n}$, i.e., there exists an invertible matrix ${\bf{Q}}$ such that ${\bf{H}}={\bf{Q}}{\bf{\Gamma}}{\bf{Q}}^{-1}$, and
$\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1.$
Therefore, we finally obtain (28). Moreover, we have
$\|{\bf{Q}}\|^{2}\leq 2\text{ and }\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}.$
Then, we prove $\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$. Since
$\mathcal{E}^{t}={\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{{\bf{R}}}^{t}\\ \end{array}\right]\text{ and }{\bf{Q}}^{-1}={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]$
taking the squared norm, we have
$\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$ $=\left\|{\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{{\bf{R}}}^{t}\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}\leq\frac{1}{4}\left\|\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}+j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{{\bf{R}}}^{t}\\ \hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}-j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{{\bf{R}}}^{t}\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}$
$\leq\|\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}\|^{2}+\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{{\bf{R}}}^{t}\|_{\mathrm{F}}^{2}.$
On the other hand, noting that
${\bf{Q}}=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}&\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}\\ -j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}&j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}\\ \end{array}\right]{\bf{Q}}_{1}^{\sf T},$
it holds that
$\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}{{\bf{R}}}^{t}\\ \end{array}\right]=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}&\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}\\ -j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}&j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}\\ \end{array}\right]{\bf{Q}}_{1}^{\sf T}\mathcal{E}^{t}=\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}({\bf{Q}}_{1,u}^{\sf T}+{\bf{Q}}_{1,l}^{\sf T})\mathcal{E}^{t}\\ -j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}({\bf{Q}}_{1,u}^{\sf T}+{\bf{Q}}_{1,l}^{\sf T})\mathcal{E}^{t}\\ \end{array}\right],$
where ${\bf{Q}}_{1,u}^{\sf T}$ and ${\bf{Q}}_{1,l}^{\sf T}$ are the upper and lower blocks of ${\bf{Q}}_{1}^{\sf T}=[{\bf{Q}}_{1,u}^{\sf T};{\bf{Q}}_{1,l}^{\sf T}]$. Then, it holds that
$\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}=\|\hat{{\bf{P}}}^{\sf T}{{\bf{X}}}^{t}\|^{2}_{\mathrm{F}}=\|\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}({\bf{Q}}_{1,u}^{\sf T}+{\bf{Q}}_{1,l}^{\sf T})\mathcal{E}^{t}\|^{2}_{\mathrm{F}}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2},$
where we used $\|{\bf{Q}}_{1,u}^{\sf T}+{\bf{Q}}_{1,l}^{\sf T}\|^{2}\leq 4$ since ${\bf{Q}}_{1}$ is a permutation matrix $\|{\bf{Q}}_{1}\|=1$. $\Box$
V Proof of Lemma 3
Proof Proof of the descent inequality (30). Since $f$ is $L$-smooth, setting $y=\bar{{\bf{x}}}^{t+1}$ and $x=\bar{{\bf{x}}}^{t}$ in (16), it gives that
$f(\bar{{\bf{x}}}^{t+1})\leq f(\bar{{\bf{x}}}^{t})+\langle\nabla f(\bar{{\bf{x}}}^{t}),\bar{{\bf{x}}}^{t+1}-\bar{{\bf{x}}}^{t}\rangle+\frac{L}{2}\|\bar{{\bf{x}}}^{t+1}-\bar{{\bf{x}}}^{t}\|^{2}.$
From (28a), i.e.,
$\bar{{\bf{x}}}^{t+1}=\bar{{\bf{x}}}^{t}-\alpha\overline{\nabla F}({\bf{X}}^{t})-\alpha\bar{{\bf{s}}}^{t},$
where $\overline{\nabla F}({\bf{X}}^{t})=\big{(}\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}({\bf{x}}^{t}_{i})\big{)}^{\sf T}$, we have
$f(\bar{{\bf{x}}}^{t+1})\leq f(\bar{{\bf{x}}}^{t})-\alpha\big{\langle}\nabla f(\bar{{\bf{x}}}^{t}),\overline{\nabla F}({\bf{X}}^{t})+\bar{{\bf{s}}}^{t}\big{\rangle}+\frac{L\alpha^{2}}{2}\|\overline{\nabla F}({\bf{X}}^{t})+\bar{{\bf{s}}}^{t}\|^{2}.$
Taking conditioned expectation with respect to $\mathcal{G}^{t}$ and by
$\mathbb{E}\!\left[\bar{{\bf{s}}}^{t}\;|\;\mathcal{G}^{t}\right]=0,\ \mathbb{E}\!\left[\|\bar{{\bf{s}}}^{t}\|^{2}\;|\;\mathcal{G}^{t}\right]\leq\frac{\sigma^{2}}{n},$
it holds that
$\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t+1})\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})\;|\;\mathcal{G}^{t}\right]-\alpha\big{\langle}\nabla f(\bar{{\bf{x}}}^{t}),\overline{\nabla F}({\bf{X}}^{t})\big{\rangle}+\mathbb{E}\!\left[\frac{L\alpha^{2}}{2}\|\overline{\nabla F}({\bf{X}}^{t})+\bar{{\bf{s}}}^{t}\|^{2}\;|\;\mathcal{F}^{t}\right]$
$=\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})\;|\;\mathcal{G}^{t}\right]-\alpha\big{\langle}\nabla f(\bar{{\bf{x}}}^{t}),\overline{\nabla F}({\bf{X}}^{t})\big{\rangle}+\frac{L\alpha^{2}}{2}\Big{(}\mathbb{E}\!\left[\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[\|\bar{{\bf{s}}}^{t}\|^{2}\;|\;\mathcal{F}^{t}\right]\Big{)}$
$\leq\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})\;|\;\mathcal{G}^{t}\right]-\alpha\big{\langle}\nabla f(\bar{{\bf{x}}}^{t}),\overline{\nabla F}({\bf{X}}^{t})\big{\rangle}+\frac{L\alpha^{2}}{2}\mathbb{E}\!\left[\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}\;|\;\mathcal{G}^{t}\right]+\frac{L\alpha^{2}\sigma^{2}}{2n}.$
Since $2\langle a,b\rangle=\|a\|^{2}+\|b\|^{2}-\|a-b\|^{2}$, we have
$-\big{\langle}\nabla f(\bar{{\bf{x}}}^{t}),\overline{\nabla F}({\bf{X}}^{t})\big{\rangle}=-\frac{1}{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}-\frac{1}{2}\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}+\frac{1}{2}\|\overline{\nabla F}({\bf{X}}^{t})-\nabla f(\bar{{\bf{x}}}^{t})\|^{2}.$
Combining the last two equations and by $\alpha L\leq\frac{1}{2}$, we get
$\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t+1})\;|\;\mathcal{G}^{t}\right]$ $\leq f(\bar{{\bf{x}}}^{t})-\frac{\alpha}{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}-\frac{\alpha}{2}\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}+\frac{\alpha}{2}\|\overline{\nabla F}({\bf{X}}^{t})-\nabla f(\bar{{\bf{x}}}^{t})\|^{2}$
$\quad+\frac{L\alpha^{2}}{2}\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n}$
$\leq f(\bar{{\bf{x}}}^{t})-\frac{\alpha}{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}-\frac{\alpha}{2}(1-\alpha L)\|\overline{\nabla F}({\bf{X}}^{t})\|^{2}+\frac{\alpha}{2}\|\overline{\nabla F}({\bf{X}}^{t})-\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n}$
$\leq f(\bar{{\bf{x}}}^{t})-\frac{\alpha}{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{\alpha}{2}\|\overline{\nabla F}({\bf{X}}^{t})-\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n}.$
By (29), i.e., $\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$, we have
$\frac{\alpha}{2}\|\overline{\nabla F}({\bf{X}}^{t})-\nabla f(\bar{{\bf{x}}}^{t})\|^{2}$ $=\frac{\alpha}{2}\|\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}(\bar{{\bf{x}}}^{t}))\|^{2}$
$\leq\frac{\alpha}{2n}\sum_{i=1}^{n}\|\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}(\bar{{\bf{x}}}^{t})\|^{2}\leq\frac{\alpha L^{2}}{2n}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq\frac{2\alpha L^{2}}{n}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}.$
Thus, the descent inequality (30) holds.
Proof of the inequality (31). Taking conditioned expectation with respect to $\mathcal{F}^{t}$, from (28f), we have
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$ $=\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|\mathbb{F}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$
$=\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right].$
Since ${\bf{E}}^{t}=\frac{(\theta_{t}-1)}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, $\mathop{\rm Prob}(\theta_{t}=1)=p$, and $\mathop{\rm Prob}(\theta_{t}=0)=1-p$, we have
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$
$=\frac{1-p}{4\chi^{2}}\Big{(}\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}+\|{\bf{Q}}^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\Big{)}\leq{\frac{2(1-p)}{\chi^{2}}}\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
Hence, it gives that
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]\leq\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}+{\frac{2(1-p)}{\chi^{2}}}\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
Taking conditioned expectation with respect to $\mathcal{G}^{t}\subset\mathcal{F}^{t}$, and using the unbiasedness of ${\bf{G}}^{t}$, we have
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{2(1-p)}{\chi^{2}}\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right].$ (50)
We first bound $\mathbb{E}\!\left[\|\mathbb{G}^{t}\|^{2}\;|\;\mathcal{G}^{t}\right]$. Recall the definition of $\mathbb{G}^{t}$.
$\mathbb{G}^{t}$ $={\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t})\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t})+\hat{{\bf{P}}}^{\sf T}(\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1}))\\ \end{array}\right]$
$={\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\underbrace{\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))+\hat{{\bf{P}}}^{\sf T}(\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1}))\\ \end{array}\right]}_{{\bf{F}}^{t}}-\alpha\underbrace{{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}\\ \end{array}\right]}_{{\bf{C}}}{\bf{S}}^{t}$
$={\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}{\bf{F}}^{t}-\alpha{\bf{C}}{\bf{S}}^{t}.$
Note that ${\bf{Q}}^{-1}={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]$ and ${\bf{I}}-\hat{{\bf{\Lambda}}}_{a}=\frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}$. It follows that
${\bf{CS}}^{t}$ $={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}\\ ({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})\hat{{\bf{P}}}^{\sf T}\\ \end{array}\right]{\bf{S}}^{t}=\frac{1}{2}{\bf{Q}}_{1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}+j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}-j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \end{array}\right],$
where ${\bf{Q}}_{1}$ is a permutation matrix $\|{\bf{Q}}_{1}\|=1$. Therefore, we have
$\|{\bf{CS}}^{t}\|^{2}_{\mathrm{F}}\leq\frac{1}{4}(\|\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}+j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\|^{2}+\|\hat{{\bf{\Lambda}}}_{a}^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}-j({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{\frac{1}{2}}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\|^{2})\leq 2\|{\bf{S}}^{t}\|^{2}_{\mathrm{F}}.$
Then, using Cauchy-Schwarz inequality, $\|\hat{{\bf{\Lambda}}}_{a}\|\leq 1$, $\|\hat{{\bf{\Lambda}}}_{b}^{2}\|\leq 2$, and $\|\hat{{\bf{P}}}^{\sf T}\|\leq 1$, we have
$\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}$ $=\|{\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}-2\alpha\langle{\bf{\Gamma}}\mathcal{E}^{t},{\bf{C}}{\bf{S}}^{t}\rangle+2\alpha^{2}\langle{\bf{Q}}^{-1}{\bf{F}}^{t},{\bf{C}}{\bf{S}}^{t}\rangle+\alpha^{2}\|{\bf{CS}}^{t}\|^{2}$
$\leq\|{\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}-2\alpha\langle{\bf{\Gamma}}\mathcal{E}^{t},{\bf{C}}{\bf{S}}^{t}\rangle+\alpha^{2}\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}+2\alpha^{2}\|{\bf{C}}{\bf{S}}^{t}\|_{\mathrm{F}}^{2}$
$\leq\|{\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}+\alpha^{2}\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}-2\alpha\langle{\bf{\Gamma}}\mathcal{E}^{t},{\bf{C}}{\bf{S}}^{t}\rangle+4\alpha^{2}\|{\bf{S}}^{t}\|_{\mathrm{F}}^{2}.$
For any matrices ${\bf{a}}$ and ${\bf{b}}$, it holds from Jensen’s inequality that $\|{\bf{a+b}}\|_{\mathrm{F}}^{2}\leq\frac{1}{\theta}\|{\bf{a}}\|_{\mathrm{F}}^{2}+\frac{1}{1-\theta}\|{\bf{b}}\|_{\mathrm{F}}^{2}$ for any $\theta\in(0,1)$. Therefore, letting $\theta=\|{\bf{\Gamma}}\|:=\gamma$, it holds that
$\|{\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\leq\frac{1}{\gamma}\|{\bf{\Gamma}}\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{1}{1-\gamma}\|\alpha{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\leq\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}}{1-\gamma}\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}.$
Since $\frac{1}{1-\gamma}>1$, we have
$\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\leq\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\alpha^{2}}{1-\gamma}\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}-2\alpha\upsilon\langle{\bf{\Gamma}}\mathcal{E}^{t},{\bf{C}}{\bf{S}}^{t}\rangle+4\alpha^{2}\|{\bf{S}}^{t}\|_{\mathrm{F}}^{2}.$
Note that
${\bf{S}}^{t}={\bf{G}}^{t}-\nabla F({\bf{X}}^{t}),\ \mathbb{E}\!\left[{\bf{S}}^{t}\;|\;\mathcal{G}^{t}\right]=0,\ \mathbb{E}\!\left[\|{\bf{S}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq n\sigma^{2}.$
It follows from this above inequality that
$\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\alpha^{2}}{1-\gamma}\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]-2\alpha\mathbb{E}\!\left[\langle{\bf{\Gamma}}\mathcal{E}^{t},{\bf{C}}{\bf{S}}^{t}\rangle\;|\;\mathcal{G}^{t}\right]+4\alpha^{2}\mathbb{E}\!\left[\|{\bf{S}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\leq\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\alpha^{2}}{1-\gamma}\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+4n\alpha^{2}\sigma^{2}.$ (51)
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ can be bounded as follows: Note that
${\bf{Q}}^{-1}={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]\text{ and }{\bf{I}}-\hat{{\bf{\Lambda}}}_{a}=\frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}.$
It follows that
${\bf{Q}}^{-1}{\bf{F}}^{t}$ $={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))\\ ({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t}))+\hat{{\bf{P}}}^{\sf T}(\nabla F(\bar{{\bf{X}}}^{t})-\nabla F(\bar{{\bf{X}}}^{t+1}))\\ \end{array}\right].$
By $\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|=\frac{2\chi}{1-\lambda_{2}}$ and $L$-smoothness of $f_{i}$, we have
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq 4L^{2}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+\frac{2\chi nL^{2}}{1-\lambda_{2}}\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t}-\bar{{\bf{x}}}^{t+1}\|^{2}\;|\;\mathcal{G}^{t}\right].$ (52)
Since $\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$, it holds that
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq 16L^{2}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\chi nL^{2}}{1-\lambda_{2}}\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t}-\bar{{\bf{x}}}^{t+1}\|^{2}\;|\;\mathcal{G}^{t}\right].$ (53)
Since $\bar{{\bf{x}}}^{t+1}=\bar{{\bf{x}}}^{t}-\alpha\overline{\nabla F}({\bf{X}}^{t})-\alpha\bar{{\bf{s}}}^{t}$, $\mathbb{E}\!\left[\bar{{\bf{s}}}^{t}\;|\;\mathcal{G}^{t}\right]=0$, and $\mathbb{E}\!\left[\|\bar{{\bf{s}}}^{t}\|^{2}\;|\;\mathcal{G}^{t}\right]\leq\frac{\sigma^{2}}{n}$, it gives that
$\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t}-\bar{{\bf{x}}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $=\mathbb{E}\!\left[\|\alpha\overline{\nabla F}({\bf{X}}^{t})+\alpha\bar{{\bf{s}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$=\alpha^{2}\mathbb{E}\!\left[\|\bar{{\bf{s}}}^{t}+(\overline{\nabla F}({\bf{X}}^{t})-\overline{\nabla F}(\bar{{\bf{X}}}^{t}))+\overline{\nabla F}(\bar{{\bf{X}}}^{t})\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\leq\alpha^{2}\mathbb{E}\!\left[\|\bar{{\bf{s}}}^{t}\|^{2}\;|\;\mathcal{G}^{t}\right]+2\alpha^{2}\|\overline{\nabla F}({\bf{X}}^{t})-\overline{\nabla F}(\bar{{\bf{X}}}^{t})\|_{\mathrm{F}}^{2}+2\alpha^{2}\|\overline{\nabla F}(\bar{{\bf{X}}}^{t})\|_{\mathrm{F}}^{2}$
$\leq\frac{\alpha^{2}\sigma^{2}}{n}+\frac{2\alpha^{2}L^{2}}{n}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+2\alpha^{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}$
$\leq\frac{\alpha^{2}\sigma^{2}}{n}+\frac{8\alpha^{2}L^{2}}{n}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+2\alpha^{2}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}.$
Then, substituting it into (53), we have
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $(16L^{2}+\frac{16\alpha^{2}L^{4}\chi}{1-\lambda_{2}})\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{2}L^{2}\chi}{1-\lambda_{2}}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{2}L^{2}\sigma^{2}\chi}{1-\lambda_{2}}.$ (54)
Thus, combining (V) and (54), it holds that
$\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=$ $\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$
$+\frac{8n\alpha^{4}L^{2}\chi}{(1-\gamma)(1-\lambda_{2})}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{4\alpha^{4}L^{2}\sigma^{2}\chi}{(1-\gamma)(1-\lambda_{2})}+4n\alpha^{2}\sigma^{2}.$ (55)
Then, we bound $\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$. Using $\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$, $\hat{{\bf{Z}}}^{t}={\bf{X}}^{t}-{\bf{R}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t}))$, and $\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}$, we have
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}({\bf{X}}^{t}-{\bf{R}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t}))\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\quad=\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}({\bf{X}}^{t}-{\bf{R}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})))\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|\alpha{\bf{S}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\quad\leq 3\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{X}}^{t}\|_{\mathrm{F}}^{2}+3\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{R}}^{t}\|_{\mathrm{F}}^{2}+3\alpha^{2}L^{2}\|{\bf{Q}}^{-1}\|^{2}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}$
$\quad\leq 3\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}.$ (56)
Combining (50) and (V), we have
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{2(1-p)}{\chi^{2}}\left(3\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}\right).$
Therefore, combining it and (V), the inequality (31) follows.
Proof of the inequality (33). Let $\bar{{\bf{e}}}^{t}\triangleq\bar{{\bf{x}}}^{t}-({\bf{x}}^{\star})^{\sf T}$. It follows from (28a), i.e., $\bar{{\bf{x}}}^{t+1}=\bar{{\bf{x}}}^{t}-\alpha\overline{\nabla F}({\bf{X}}^{t})-\alpha\bar{{\bf{s}}}^{t}$, Assumption 4, and $\sum_{i=1}^{n}\nabla f_{i}({\bf{x}}^{\star})=0$ that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{e}}}^{t+1}\big{\|}^{2}\;|\;\mathcal{G}^{t}\right]=\big{\|}\bar{{\bf{e}}}^{t}-\frac{\alpha}{n}\sum_{i=1}^{n}(\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}({\bf{x}}^{\star}))\big{\|}^{2}+\alpha^{2}\mathbb{E}\!\left[\|\bar{{\bf{s}}}^{k}\|^{2}\;|\;\mathcal{G}^{t}\right]$
$\leq\big{\|}\bar{{\bf{e}}}^{t}-\frac{\alpha}{n}\sum_{i=1}^{n}(\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}({\bf{x}}^{\star}))\big{\|}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}$
$=\|\bar{{\bf{e}}}^{t}\|^{2}+\alpha^{2}\Big{\|}\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}({\bf{x}}^{\star}))\Big{\|}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-\frac{2\alpha}{n}\sum_{i=1}^{n}\langle\nabla f_{i}({\bf{x}}_{i}^{t}),\bar{{\bf{e}}}^{t}\rangle.$ (57)
It follows from the $L$-smoothness of $f$ and $f_{i}$ and Jensen’s inequality that
$\alpha^{2}\Big{\|}\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}({\bf{x}}^{\star}))\Big{\|}^{2}=\alpha^{2}\Big{\|}\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}(\bar{{\bf{x}}}^{t})+\nabla f_{i}(\bar{{\bf{x}}}^{t})-\nabla f_{i}({\bf{x}}^{\star}))\Big{\|}^{2}$
$\leq 2\alpha^{2}\Big{\|}\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}(\bar{{\bf{x}}}^{t}))\Big{\|}^{2}+2\alpha^{2}\Big{\|}\frac{1}{n}\sum_{i=1}^{n}(\nabla f_{i}(\bar{{\bf{x}}}^{t})-\nabla f_{i}({\bf{x}}^{\star}))\Big{\|}^{2}$
$\leq\frac{2\alpha^{2}}{n}\sum_{i=1}^{n}\Big{\|}\nabla f_{i}({\bf{x}}_{i}^{t})-\nabla f_{i}(\bar{{\bf{x}}}^{t})\Big{\|}^{2}+2\alpha^{2}\Big{\|}\nabla f_{i}(\bar{{\bf{x}}}^{t})-\nabla f_{i}({\bf{x}}^{\star})\Big{\|}^{2}$
$\leq\frac{2\alpha^{2}L^{2}}{n}\|{\bf{X}}^{t}-{\bf{1}}\bar{{\bf{x}}}^{t}\|_{\mathrm{F}}^{2}+4\alpha^{2}L\big{(}f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})-\langle\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star},\nabla f({\bf{x}}^{\star})\rangle\big{)}$
$=\frac{2\alpha^{2}L^{2}}{n}\|{\bf{X}}^{t}-{\bf{1}}\bar{{\bf{x}}}^{t}\|_{\mathrm{F}}^{2}+4\alpha^{2}L(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})).$ (58)
Then, we consider the bound of $-\frac{2\alpha}{n}\sum_{i=1}^{n}\langle\nabla f_{i}({\bf{x}}_{i}^{t}),\bar{{\bf{e}}}^{t}\rangle$. Since $f_{i}$ is $L$-smooth and $\mu$-strongly convex, by (14), it gives that
$-\frac{2\alpha}{n}\sum_{i=1}^{n}\langle\nabla f_{i}({\bf{x}}_{i}^{t}),\bar{{\bf{e}}}^{t}\rangle=\frac{2\alpha}{n}\sum_{i=1}^{n}\big{(}-\langle\nabla f_{i}({\bf{x}}_{i}^{t}),\bar{{\bf{x}}}^{t}-{\bf{x}}_{i}^{t}\rangle-\langle\nabla f_{i}({\bf{x}}_{i}^{t}),{\bf{x}}^{t}_{i}-{\bf{x}}^{\star}\rangle\big{)}$
$\leq\frac{2\alpha}{n}\sum_{i=1}^{n}\Big{(}-f_{i}(\bar{{\bf{x}}}^{t})+f_{i}({\bf{x}}_{i}^{t})+\frac{L}{2}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{t}_{i}\|^{2}-\frac{\mu}{2}\|{\bf{x}}_{i}^{t}-{\bf{x}}^{\star}\|^{2}-f_{i}({\bf{x}}_{i}^{t})+f_{i}({\bf{x}}^{\star})\Big{)}$
$\leq-2\alpha(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star}))+\frac{\alpha L}{n}\sum_{i=1}^{n}\|\bar{{\bf{x}}}^{t}-{\bf{x}}_{i}^{t}\|^{2}-\mu\alpha\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}$
$=-2\alpha(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star}))+\frac{\alpha L}{n}\|{\bf{X}}^{t}-{\bf{1}}\bar{{\bf{x}}}^{t}\|_{\mathrm{F}}^{2}-\mu\alpha\|\bar{{\bf{e}}}^{t}\|^{2}.$ (59)
Substituting (V) and (V) into (V), and using $f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})\geq 0$, we have
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{e}}}^{t+1}\big{\|}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $(1-\mu\alpha)\|\bar{{\bf{e}}}^{t}\|^{2}+\Big{(}\frac{\alpha L}{n}+\frac{2\alpha^{2}L^{2}}{n}\Big{)}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-2\alpha(1-2\alpha L)(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})).$ (60)
Since $\alpha\leq\frac{1}{4L}$, it holds that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $(1-\mu\alpha)\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\Big{(}\frac{\alpha L}{n}+\frac{2\alpha^{2}L^{2}}{n}\Big{)}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-2\alpha(1-2\alpha L)(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star}))$
$\leq$ $(1-\mu\alpha)\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{3\alpha L}{2n}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-\alpha(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})).$
Combining with $\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$, we complete the proof. $\Box$
VI Proof of Theorem 3
Proof From the condition of stepsize, we have
$\alpha\leq\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L}\Longrightarrow\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\leq 3.$
Then, it follows from the definition of $\tilde{\gamma}$ (32) that
$\tilde{\gamma}$ $=\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2(1-p)\big{(}3+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\big{)}}{\chi^{2}}$
$\leq\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{12(1-p)}{\chi^{2}}$
$\leq\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{12(1-p)}{\chi^{2}}.$
To ensure $\tilde{\gamma}\leq\frac{1+\gamma}{2}$, we need to choose $\alpha$ and $\chi$ such that
$\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{12(1-p)}{\chi^{2}}\leq\frac{1-\gamma}{2}.$
By solving these inequalities
$\frac{32\alpha^{2}L^{2}}{1-\gamma}\leq\frac{1-\gamma}{6},\quad\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\leq\frac{1-\gamma}{6},\quad\frac{12(1-p)}{\chi^{2}}\leq\frac{1-\gamma}{6},$
and using $\gamma=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}$, we have
$\alpha\leq\min\left\{\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}.$
Thus, it implies that if the condition of $\alpha$ and $\chi$ in this Lemma holds, then $\tilde{\gamma}\leq\frac{1+\gamma}{2}<1$.
Define the Lyapunov function
$\mathcal{L}^{t}=f(\bar{{\bf{x}}}^{t})-f^{\star}+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|^{2}.$
Since $\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}}$ and $\frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma}$, we have
$\frac{\frac{2\chi}{1-\lambda_{2}}}{(1-\gamma)^{2}}\leq\frac{32\chi^{3}}{(1-\lambda_{2})^{3}}\text{ and }\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\leq\frac{1024\alpha^{4}L^{4}\chi^{3}}{(1-\lambda_{2})^{3}}.$
Thus, it gives that
$\alpha\leq\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{8\chi^{3}}}\frac{1}{4L}\Rightarrow\frac{1}{2}<1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}.$
Then, since $\alpha\leq\frac{1}{2L}$, by (30) and (31), it gives
$\mathbb{E}\!\left[\mathcal{L}^{t+1}\;|\;\mathcal{G}^{t}\right]\leq$ $f(\bar{{\bf{x}}}^{t})-f^{\star}-\frac{\alpha}{2}\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}+\frac{2\alpha L^{2}}{n}\big{\|}\mathcal{E}^{t}\big{\|}_{\mathrm{F}}^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n}$
$+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\Big{(}\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}}\Big{)}$
$=$ $f(\bar{{\bf{x}}}^{t})-f^{\star}+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}-\frac{\alpha}{2}\Big{(}1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\Big{)}\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}$
$+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}$
$\leq$ $\mathcal{L}^{t}-\frac{\alpha}{4}\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}},$
where the last inequality holds because the condition (34) implies $\frac{1}{2}<1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}$. Taking full expectation, we have
$\mathbb{E}\!\left[\mathcal{L}^{t+1}\right]\leq\mathbb{E}\!\left[\mathcal{L}^{t}\right]-\frac{\alpha}{4}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
Summing it over $t=0,1,\cdots,T-1$, we can obtain
$\frac{\alpha}{4}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]\leq\mathcal{L}^{0}+T\Big{(}\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}\Big{)},$
which implies that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]\leq\frac{4\mathcal{L}^{0}}{\alpha T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{16\sigma^{2}L^{4}\alpha^{4}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{16L^{2}\sigma^{2}\alpha^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
Since ${\bf{X}}^{0}=[{\bf{x}}^{0},\cdots,{\bf{x}}^{0}]^{\sf T}$, by [42, (75)], we have $\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq 2\alpha^{2}\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|\|\nabla F({\bf{X}}^{0})-{\bf{1}}_{n}\otimes(\nabla f({\bf{x}}^{0}))^{\sf T}\|^{2}$. Notice that $\varsigma^{2}_{0}=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\bar{{\bf{x}}}^{0})-\nabla f(\bar{{\bf{x}}}^{0})\|^{2}$. It holds that
$\mathcal{L}^{0}$ $=f(\bar{{\bf{x}}}^{0})-f^{\star}+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}$
$=f(\bar{{\bf{x}}}^{0})-f^{\star}+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\Big{(}2\alpha^{2}\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|\|\nabla F({\bf{X}}^{0})-{\bf{1}}_{n}\otimes(\nabla f({\bf{x}}^{0}))^{\sf T}\|^{2}\Big{)}$
$\leq f(\bar{{\bf{x}}}^{0})-f^{\star}+\frac{32\chi^{2}\alpha^{3}L^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$ (61)
Using (61) and
$\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}},\ \frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma},$
we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]$ $\leq\frac{4(f(\bar{{\bf{x}}}^{0})-f^{*})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{1024\sigma^{2}L^{4}\alpha^{4}\chi^{3}}{n(1-\lambda_{2})^{3}}+\frac{128\chi\alpha^{2}L^{2}\sigma^{2}(2\chi^{2}+(1-p))}{(1-\lambda_{2})\chi^{2}}.$
Since $\alpha\leq\frac{1-\lambda_{2}}{32\sqrt{3}\chi L}$, we have $\frac{1024\sigma^{2}L^{4}\alpha^{4}\chi^{3}}{n(1-\lambda_{2})^{3}}\leq\frac{\alpha^{2}L^{2}\sigma^{2}\chi}{3n(1-\lambda_{2})}\leq\frac{\alpha^{2}L^{2}\sigma^{2}\chi}{2(1-\lambda_{2})}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]$ $\leq\frac{4(f(\bar{{\bf{x}}}^{0})-f^{*})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}\chi^{3}+256\chi\alpha^{2}L^{2}\sigma^{2}(2\chi^{2}+(1-p))}{2(1-\lambda_{2})\chi^{2}},$
i.e, (35) holds. $\Box$
VII Proof of Corollary 2
Proof We derive a tighter rate by carefully selecting the step size similar to [31]. We rewrite (35) as
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]\leq\underbrace{\frac{c_{0}}{\alpha T}+c_{1}\alpha+c_{2}\alpha^{2}}_{:=\Psi_{T}}+\frac{a_{0}\alpha^{2}}{T},$ (62)
where
$c_{0}=4(f(\bar{{\bf{x}}}^{0})-f^{*}),\ c_{1}=\frac{2L\sigma^{2}}{n},\ c_{2}=\frac{L^{2}\sigma^{2}\big{(}\chi^{3}+256\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}},\ a_{0}=\frac{128\chi^{2}L^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$ (63)
From the condition of stepsize, we have
$\alpha\leq\frac{1}{\underline{\alpha}}=\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\}=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right).$
Setting
$\alpha=\min\left\{\left(\frac{c_{0}}{c_{1}T}\right)^{\frac{1}{2}},\left(\frac{c_{0}}{c_{2}T}\right)^{\frac{1}{3}},\frac{1}{\underline{\alpha}}\right\},$
we have the following cases.
- When $\alpha=\frac{1}{\underline{\alpha}}$ and is smaller than both $\left(\frac{c_{0}}{c_{1}T}\right)^{\frac{1}{2}}$ and $\left(\frac{c_{0}}{c_{2}T}\right)^{\frac{1}{3}}$, then
$\Psi_{T}=\frac{c_{0}}{\alpha T}+c_{1}\alpha+c_{2}\alpha^{2}=\frac{\alpha c_{0}}{T}+\frac{c_{1}}{\underline{\alpha}}+\frac{c_{2}}{\underline{\alpha}^{2}}\leq\frac{\alpha c_{0}}{T}+c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{T}\right)^{\frac{1}{2}}+c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{T}\right)^{\frac{2}{3}}.$
- When $\alpha=\left(\frac{c_{0}}{c_{1}T}\right)^{\frac{1}{2}}\leq\left(\frac{c_{0}}{c_{2}T}\right)^{\frac{1}{3}}$, then
$\Psi_{T}\leq 2c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{T}\right)^{\frac{1}{2}}+c_{2}\left(\frac{c_{0}}{c_{1}T}\right)\leq 2c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{T}\right)^{\frac{1}{2}}+c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{T}\right)^{\frac{2}{3}}.$
- When $\alpha=\left(\frac{c_{0}}{c_{2}T}\right)^{\frac{1}{3}}\leq\left(\frac{c_{0}}{c_{1}T}\right)^{\frac{1}{2}}$, then
$\Psi_{T}\leq 2c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{T}\right)^{\frac{2}{3}}+c_{1}\left(\frac{c_{0}}{c_{2}T}\right)^{\frac{1}{3}}\leq 2c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{T}\right)^{\frac{2}{3}}+c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{T}\right)^{\frac{1}{2}}.$
Combining the above three cases together it holds that
$\Psi_{T}=\frac{c_{0}}{\alpha T}+c_{1}\alpha+c_{2}\alpha^{2}\leq 2c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{T}\right)^{\frac{1}{2}}+2c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{T}\right)^{\frac{2}{3}}+\frac{\alpha c_{0}}{T}.$
Substituting the above into (62), we conclude that
$\frac{1}{R}\sum_{r=0}^{R-1}\mathcal{E}_{r}\leq 2c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{R}\right)^{\frac{1}{2}}+2c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{R}\right)^{\frac{2}{3}}+\frac{\left(\underline{\alpha}c_{0}+a_{0}/\underline{\alpha}^{2}\right)}{R}$ (64)
Therefore, from (64) and plugging the parameters (63)
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\mathcal{O}\left(\sqrt{\frac{L(f(\bar{{\bf{x}}}^{0})-f^{\star})\sigma^{2}}{nT}}\right)+\mathcal{O}\left(\sqrt[3]{\frac{\chi^{3}+\chi(1-p)}{(1-\lambda_{2})\chi^{2}}}\left(\frac{L(f(\bar{{\bf{x}}}^{0})-f^{\star})\sigma}{T}\right)^{\frac{2}{3}}\right)+\mathcal{O}\left(\frac{\frac{\chi L(f(\bar{{\bf{x}}}^{0})-f^{\star})}{1-\lambda_{2}}+\varsigma_{0}^{2}}{T}\right),$
i.e., the rate (36) holds. $\Box$
VIII Proof of Theorem 4
Proof Plugging $\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}\leq 2L(f(\bar{{\bf{x}}}^{t})-f^{\star})$ into (31) gives
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{8n\alpha^{4}L^{3}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}}.$ (65)
Similar as Lemma 3, we know that
$\alpha\leq\min\left\{\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}\leq\frac{1+\gamma}{2}<1.$
Define the Lyapunov function
$\mathcal{L}_{\mathrm{c}}^{t}=\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}.$
Since $\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}}$ and $\frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma}$, we have
$\frac{\frac{2\chi}{1-\lambda_{2}}}{(1-\gamma)^{2}}\leq\frac{32\chi^{3}}{(1-\lambda_{2})^{3}}\text{ and }\frac{24\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\leq\frac{96\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\gamma)^{2}}.$
It gives that
$\alpha\leq\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\Rightarrow\frac{1}{2}<1-\frac{48\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}.$
Thus, according to (33), (65), and $\mu=0$, we have
$\mathbb{E}\!\left[\mathcal{L}_{\mathrm{c}}^{t+1}\;|\;\mathcal{G}^{t}\right]\leq$ $\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-\alpha(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star}))$
$+\frac{6\alpha L}{n(1-\tilde{\gamma})}\Big{(}\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{8n\alpha^{4}L^{3}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}}\Big{)}$
$=$ $\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}-\alpha\Big{(}1-\frac{48\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\Big{)}(f(\bar{{\bf{x}}}^{t})-f^{\star})$
$+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}$
$\leq$ $\mathcal{L}_{\mathrm{c}}^{t}-\frac{\alpha}{2}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
Taking full expectation, we have
$\mathbb{E}\!\left[\mathcal{L}_{\mathrm{c}}^{t+1}\right]\leq\mathbb{E}\!\left[\mathcal{L}_{\mathrm{c}}^{t}\right]-\frac{\alpha}{2}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$ (66)
Summing the inequality (66) over $t=0,1,\cdots,T-1$, we can obtain
$\frac{\alpha}{2}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{L}_{\mathrm{c}}^{0}+T\Big{(}+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}\Big{)},$
which implies that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\frac{2\mathcal{L}_{\mathrm{c}}^{0}}{\alpha T}+\frac{2\alpha\sigma^{2}}{n}+\frac{24\alpha^{4}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{24\alpha^{2}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$ (67)
Since ${\bf{X}}^{0}=[{\bf{x}}^{0},\cdots,{\bf{x}}^{0}]^{\sf T}$, similar as (61), we have
$\mathcal{L}_{\mathrm{c}}^{0}$ $=\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n(1-\tilde{\gamma})}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}+\frac{96\chi^{2}\alpha^{3}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$ (68)
Substituting (68) into (67) and using
$\tilde{\gamma}\leq\frac{1+\gamma}{2}<1,\ \frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}},$
we can derive that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{1536\chi^{3}\alpha^{4}L^{3}\sigma^{2}}{n(1-\lambda_{2})^{3}}+\frac{192\alpha^{2}L\sigma^{2}\chi(2\chi^{2}+(1-p))}{(1-\lambda_{2})\chi^{2}}.$
Since $\alpha\leq\frac{1-\lambda_{2}}{32\sqrt{3}\chi L}$, we have $\frac{1536\sigma^{2}L^{3}\alpha^{4}\chi^{3}}{n(1-\lambda_{2})^{3}}\leq\frac{\alpha^{2}L\sigma^{2}\chi}{2n(1-\lambda_{2})}\leq\frac{\alpha^{2}L\sigma^{2}\chi}{2(1-\lambda_{2})}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L\sigma^{2}\chi^{3}+384\alpha^{2}L\sigma^{2}\chi(2\chi^{2}+(1-p))}{2(1-\lambda_{2})\chi^{2}}.$
i.e., (38) holds. $\Box$
IX Proof of Corollary 3
Proof Then, we derive a tighter rate by carefully selecting the step size similar to Corollary 2. From the condition of stepsize, we have
$\alpha\leq\frac{1}{\underline{\alpha}}=\min\left\{\frac{1}{4L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\right\}=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right).$
Similar as the proof of Theorem 3, it follows that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\underbrace{\frac{c_{0}}{\alpha T}+c_{1}\alpha+c_{2}\alpha^{2}}_{:=\Psi_{T}}+\frac{a_{0}\alpha^{2}}{T},$
where
$c_{0}=2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2},\ c_{1}=\frac{2\sigma^{2}}{n},\ c_{2}=\frac{L\sigma^{2}\big{(}\chi^{3}+384\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}},\ a_{0}=\frac{192\chi^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$
Then, the following rate can be obtained by following the same arguments used for the noncovex case,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\mathcal{O}\left(\sqrt{\frac{\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}\sigma^{2}}{nT}}+\sqrt[3]{\frac{\chi^{3}+\chi(1-p)}{(1-\lambda_{2})\chi^{2}}}L^{\frac{1}{3}}\Big{(}\frac{\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}\sigma}{T}\Big{)}^{\frac{2}{3}}+\frac{\frac{\chi L\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{1-\lambda_{2}}+\varsigma_{0}^{2}}{T}\right),$
i.e., (39) holds. $\Box$
X Proof of Theorem 5
Proof Recall (31)
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\;|\;\mathcal{G}^{t}\right]\leq(1-\mu\alpha)\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-\alpha(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})),$
From (31) and (33), we have
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\;|\;\mathcal{G}^{t}\right]\leq(1-\mu\alpha)\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n},$
and
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}},$
where the last inequality follows from $\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}\leq L^{2}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}$. Similar as Lemma 3, we know that
$\alpha\leq\min\left\{\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}\leq\frac{1+\gamma}{2}<1.$
Since $\alpha\leq\frac{1-\lambda_{2}}{32\sqrt{3}\chi L}$ and $\frac{\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\leq\frac{8\chi^{2}}{(1-\lambda_{2})^{2}}$, we have $\frac{\alpha^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\leq\frac{1}{384L^{2}}$. Thus, it holds that
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{n\alpha^{2}L^{2}}{96}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{n\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}.$
Then, it follows that
$\left[\begin{array}[]{c}\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]\leq\underbrace{\left[\begin{array}[]{cc}1-\mu\alpha&6\alpha L\\ \frac{\alpha^{2}L^{2}}{96}&\frac{1+\gamma}{2}\\ \end{array}\right]}_{:=A}\left[\begin{array}[]{c}\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big{\|}^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]+\underbrace{\left[\begin{array}[]{c}\frac{\alpha^{2}\sigma^{2}}{n}\\ \frac{\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \end{array}\right]}_{:=b}.$ (77)
Note that
$\alpha\leq\min\left\{\frac{72\mu}{L^{2}},\frac{1-\gamma}{12L+\nicefrac{{\mu}}{{2}}}\right\}\Longrightarrow\|A\|\leq\|A\|_{1}=\max\left\{1-\mu\alpha+\frac{\alpha^{2}L^{2}}{96},6\alpha L+\frac{1+\gamma}{2}\right\}\leq 1-\frac{\mu\alpha}{4}<1.$
Since $\|A\|<1$, we can iterate inequality (77) to get
$\left[\begin{array}[]{c}\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]\leq A^{t}\left[\begin{array}[]{c}\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]+\sum_{\ell=0}^{t-1}A^{\ell}b\leq A^{t}\left[\begin{array}[]{c}\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]+(I-A)^{-1}b.$
Taking the 1-induced-norm and using properties of the (induced) norms, it holds that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big{\|}^{2}\right]+\frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}\right]\leq\|A^{t}\|_{1}a_{0}+\|(I-A)^{-1}b\|_{1}\leq\|A\|_{1}^{t}a_{0}+\|(I-A)^{-1}b\|_{1},$ (78)
where $a_{0}=\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{1}{n}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}$. We now bound the last term by noting that
$(I-A)^{-1}b$ $=\left[\begin{array}[]{cc}\mu\alpha&-6\alpha L\\ -\frac{\alpha^{2}L^{2}}{96}&\frac{1-\gamma}{2}\\ \end{array}\right]^{-1}b=\frac{1}{\mathrm{det}(I-A)}\left[\begin{array}[]{cc}\frac{1-\gamma}{2}&6\alpha L\\ \frac{\alpha^{2}L^{2}}{96}&\mu\alpha\\ \end{array}\right]b$
$=\frac{1}{\mu\alpha(1-\gamma)(\frac{1}{2}-\frac{\alpha^{3}L^{3}}{16\mu(1-\gamma)})}\left[\begin{array}[]{cc}\frac{1-\gamma}{2}&6\alpha L\\ \frac{\alpha^{2}L^{2}}{96}&\mu\alpha\\ \end{array}\right]\left[\begin{array}[]{c}\frac{\alpha^{2}\sigma^{2}}{n}\\ \frac{\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \end{array}\right]$
$\leq\frac{4}{\alpha\mu(1-\gamma)}\left[\begin{array}[]{c}\frac{(1-\gamma)\alpha^{2}\sigma^{2}}{2n}+\frac{6L\alpha^{3}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \frac{\alpha^{4}L^{2}\sigma^{2}}{96n}+\frac{\mu\alpha^{3}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \end{array}\right],$
where the last step holds for $\alpha\leq\sqrt[3]{4\mu(1-\gamma)}\frac{1}{L}$. Therefore,
$\|(I-A)^{-1}b\|_{1}\leq\frac{2\alpha\sigma^{2}}{\mu n}+\frac{(6L\alpha^{2}\sigma^{2}+\mu\alpha^{2}\sigma^{2})(192\chi^{2}+(4\chi^{2}+2(1-p)))}{48\mu(1-\gamma)\chi^{2}}.$
Substituting the above into (78) and using $\|A\|^{t}_{1}\leq(1-\frac{\alpha\mu}{4})^{t}$ and $\mu\leq L$, we obtain
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq\Big{(}1-\frac{\alpha\mu}{4}\Big{)}^{t}a_{0}+\frac{2\alpha\sigma^{2}}{\mu n}+\frac{7L\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{48\mu(1-\gamma)\chi^{2}}.$
Since ${\bf{X}}^{0}=[{\bf{x}}^{0},\cdots,{\bf{x}}^{0}]^{\sf T}$, by [42, (75)], we have $\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq 2\alpha^{2}\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|\|\nabla F({\bf{X}}^{0})-{\bf{1}}_{n}\otimes(\nabla f({\bf{x}}^{0}))^{\sf T}\|^{2}=\frac{2n\alpha^{2}\varsigma_{0}^{2}}{1-\gamma}$. Note that $\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}}$. It holds that
$a_{0}=\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{1}{n}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{1-\lambda_{2}}$
Thus, we finally obtain (41). $\Box$
XI Proof of Corollary 4
Proof Recall (41)
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\big{\|}^{2}\right]$ $\leq\Big{(}1-\frac{\alpha\mu}{4}\Big{)}^{T}\Big{(}c_{0}+b_{0}\alpha^{2}\Big{)}+c_{1}\alpha+c_{2}\alpha^{2}$
$\leq{\mathrm{exp}}\Big{(}-\frac{\alpha\mu}{2}T\Big{)}\Big{(}c_{0}+b_{0}\alpha^{2}\Big{)}+c_{1}\alpha+c_{2}\alpha^{2},$ (79)
where
$c_{0}=\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2},\ b_{0}=\frac{8\chi\varsigma_{0}^{2}}{1-\lambda_{2}},\ c_{1}=\frac{2\sigma^{2}}{\mu n},\ c_{2}=\frac{7L\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{12\mu(1-\lambda_{2})\chi}.$
From the setpsize condition, we have
$\alpha\leq\frac{1}{\underline{\alpha}}\triangleq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\frac{72\mu}{L^{2}},\frac{1-\gamma}{12L+\nicefrac{{\mu}}{{2}}},\sqrt[3]{4\mu(1-\gamma)}\frac{1}{L}\right\}=\mathcal{O}\left(\frac{\mu(1-\lambda_{2})}{\chi L^{2}}\right).$
Now we select $\alpha=\min\left\{\frac{\ln\left(\max\left\{1,\mu\left(c_{0}+b_{0}/\underline{\alpha}^{2}\right)T/c_{1}\right\}\right)}{\mu T},\frac{1}{\underline{\alpha}}\right\}\leq\frac{1}{\underline{\alpha}}$ to get the following cases.
- If $\alpha=\frac{\ln\left(\max\left\{1,\mu\left(c_{0}+b_{0}/\underline{\alpha}^{2}\right)T/c_{1}\right\}\right)}{\mu T}\leq\frac{1}{\underline{\alpha}}$ then
$\exp\left(-\frac{\alpha\mu}{2}T\right)\left(c_{0}+\alpha^{2}b_{0}\right)$ $\leq\tilde{\mathcal{O}}\left(\left(c_{0}+\frac{b_{0}}{\underline{\alpha}^{2}}\right)\exp\left[-\ln\left(\max\left\{1,\mu\left(c_{0}+\frac{b_{0}}{\underline{\alpha}^{2}}\right)T/c_{1}\right\}\right)\right]\right)=\mathcal{O}\left(\frac{c_{1}}{\mu T}\right)$
- Otherwise $\alpha=\frac{1}{\underline{\alpha}}\leq\frac{\ln\left(\max\left\{1,\mu\left(c_{0}+b_{0}/\underline{\alpha}^{2}\right)/c_{1}\right\}\right)}{\mu T}$ and
$\exp\left(-\frac{\alpha\mu}{2}T\right)\left(c_{0}+\alpha^{2}b_{0}\right)=\tilde{\mathcal{O}}\left(\exp\left[-\frac{\mu T}{2\underline{\alpha}}\right]\left(c_{0}+\frac{b_{0}}{\underline{\alpha}^{2}}\right)\right).$
Collecting these cases together into (XI), we have
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\big{\|}^{2}\right]$ $\leq\exp\left(-\frac{\alpha\mu}{2}T\right)\left(c_{0}+\alpha^{2}b_{0}\right)+c_{1}\alpha+c_{2}\alpha^{2}$
$\leq\tilde{\mathcal{O}}\left(\frac{c_{1}}{\mu T}\right)+\tilde{\mathcal{O}}\left(\frac{c_{2}}{\mu^{2}T^{2}}\right)+\tilde{\mathcal{O}}\left(\exp\left[-\frac{\mu T}{2\underline{\alpha}}\right]\left(c_{0}+\frac{b_{0}}{\underline{\alpha}^{2}}\right)\right).$
Therefore, (42) holds. $\Box$
XII Proof of Lemma 5
Proof Note that ProxSkip (43) has the following equivalently updates
$\widetilde{{\bf{Z}}}^{t}$ $=\widetilde{{\bf{X}}}^{t}-{\bf{W}}_{b}\widetilde{{\bf{U}}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t}),$ (80a)
$\widetilde{{\bf{X}}}^{t+1}$ $={\bf{W}}_{a}\widetilde{{\bf{Z}}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t},$ (80b)
$\widetilde{{\bf{U}}}^{t+1}$ $=\widetilde{{\bf{U}}}^{t}+\frac{p}{2\chi}{\bf{W}}_{b}\widetilde{{\bf{Z}}}^{t}+p{\bf{E}}^{t}.$ (80c)
We rewrite the recursion (80) into the following matrix representation:
$\left[\begin{array}[]{c}\widetilde{{\bf{X}}}^{t+1}\\ \widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]=\left[\begin{array}[]{cc}{\bf{W}}_{a}&-{\bf{W}}_{a}{\bf{W}}_{b}\\ \frac{p}{2\chi}{\bf{W}}_{b}&{\bf{I}}-\frac{p}{2\chi}{\bf{W}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}\widetilde{{\bf{X}}}^{t}\\ \widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}{\bf{W}}_{a}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}{\bf{W}}_{b}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-{\bf{W}}_{b}{\bf{E}}^{t}\\ p{\bf{E}}^{t}\\ \end{array}\right].$
Multiplying both sides of the above by $\mathrm{diag}\{{\bf{P}}^{-1},{\bf{P}}^{-1}\}$ on the left and using (27), we have
$\left[\begin{array}[]{c}{\bf{P}}^{-1}\widetilde{{\bf{X}}}^{t+1}\\ {\bf{P}}^{-1}\widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}{\bf{P}}^{-1}\widetilde{{\bf{X}}}^{t}\\ {\bf{P}}^{-1}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}{\bf{P}}^{-1}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}{\bf{P}}^{-1}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}{\bf{P}}^{-1}{\bf{E}}^{t}\\ p{\bf{P}}^{-1}{\bf{E}}^{t}\\ \end{array}\right].$
Since $\widetilde{{\bf{U}}}^{t}$ lies in the range space of ${\bf{W}}_{b}$, we have ${\bf{1}}^{\sf T}\widetilde{{\bf{U}}}^{t}=0,\ t\geq 0$. By the structure of ${\bf{P}}$, we have
${\bf{P}}^{-1}\widetilde{{\bf{X}}}^{t}=\left[\begin{array}[]{c}\bar{{\bf{e}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\\ \end{array}\right],\ {\bf{P}}^{-1}\widetilde{{\bf{U}}}^{t}=\left[\begin{array}[]{c}0\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right],\ {\bf{P}}^{-1}\nabla F({\bf{X}}^{t})=\left[\begin{array}[]{c}\overline{\nabla F}({\bf{X}}^{t})\\ \hat{{\bf{P}}}^{\sf T}\nabla F({\bf{X}}^{t})\\ \end{array}\right],\ {\bf{P}}^{-1}{\bf{E}}^{t}=\left[\begin{array}[]{c}0\\ \hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
Therefor, it holds that
$\bar{{\bf{e}}}^{t+1}$ $=\bar{{\bf{e}}}^{t}-\alpha\overline{\nabla F}({\bf{X}}^{t})-\alpha\bar{{\bf{s}}}^{t},$
$\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t+1}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]$ $=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ p\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
Let
${\bf{H}}^{\mathrm{s}}=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]=\left[\begin{array}[]{cc}{\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})&-({\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}}))\sqrt{{\bf{I}}-\hat{{\bf{\Lambda}}}}\\ \frac{p}{2\chi}\sqrt{{\bf{I}}-\hat{{\bf{\Lambda}}}}&{\bf{I}}-\frac{p}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})\\ \end{array}\right]$
where $\hat{{\bf{\Lambda}}}=\mathrm{diag}\{\lambda_{2},\cdots,\lambda_{n}\}$, and $\lambda_{i}\in(-1,1)$. Since the blocks of ${\bf{H}}^{\mathrm{s}}$ are diagonal matrices, there exists a permutation matrix ${\bf{Q}}^{\mathrm{s}}_{1}$ such that ${\bf{Q}}^{\mathrm{s}}_{1}{\bf{H}}^{\mathrm{s}}({\bf{Q}}^{\mathrm{s}}_{1})^{\sf T}=\mathrm{blkdiag}\{H^{\mathrm{s}}_{i}\}_{i=2}^{n}$, where
$H^{\mathrm{s}}_{i}=\left[\begin{array}[]{cc}1-\frac{1}{2\chi}(1-\lambda_{i})&-(1-\frac{1}{2\chi}(1-\lambda_{i}))\sqrt{1-\lambda_{i}}\\ \frac{p}{2\chi}\sqrt{1-\lambda_{i}}&1-\frac{p}{2\chi}(1-\lambda_{i})\\ \end{array}\right].$
Setting $\nu_{i}=1-\frac{1}{2\chi}(1-\lambda_{i})$, we have $\nu_{i}\in(0,1)$ and $H_{i}$ can be rewritten as
$H^{\mathrm{s}}_{i}=\left[\begin{array}[]{cc}\nu_{i}&-\nu_{i}\sqrt{2\chi(1-\nu_{i})}\\ \frac{p}{2\chi}\sqrt{2\chi(1-\nu_{i})}&1-p(1-\nu_{i})\\ \end{array}\right].$
Since
$\mathrm{Tr}(H^{\mathrm{s}}_{i})=(1+p)\nu_{i}+(1-p),\quad\mathrm{det}(H^{\mathrm{s}}_{i})=\nu_{i},$
the eigenvalues of $H_{i}$ are
$\gamma_{(1,2),i}$ $=\frac{1}{2}\Big{[}\mathrm{Tr}(H^{\mathrm{s}}_{i})\pm\sqrt{\mathrm{Tr}(H^{\mathrm{s}}_{i})^{2}-4\mathrm{det}(H^{\mathrm{s}}_{i})}\Big{]}$
$=\frac{1}{2}\Big{[}(1+p)\nu_{i}+(1-p)\Big{]}\pm\frac{1}{2}\sqrt{\underbrace{(1+p)^{2}\nu_{i}^{2}+(2(1+p)(1-p)-4)\nu_{i}+(1-p)^{2}}_{:=\Delta_{i}(\nu_{i},p)}}.$
Consider the sign of $\Delta_{i}(\nu_{i},p)$. Note that $\Delta_{i}(\nu_{i},p)$ is a quadratic function on $\nu_{i}$, and
$(1+p)^{2}>0,\ \Delta_{i}(0,p)=(1-p)^{2},\ \Delta_{i}(1,p)=0,\ \Delta_{i}(c_{i},p)=0,\text{ where }c_{i}=\frac{(1-p)^{2}}{(1+p)^{2}}<1.$
We have
$\left\{\begin{array}[]{cc}\Delta_{i}(\nu_{i},p)>0,&\nu_{i}\in(0,c_{i})\\ \Delta_{i}(\nu_{i},p)<0,&\nu_{i}\in(c_{i},1)\end{array}\right..$
Since $\nu_{i}=1-\frac{1}{2\chi}(1-\lambda_{i})\geq 1-\frac{1}{2\chi}(1-\lambda_{n}),i=2,\ldots,n$ and $\lambda_{n}\in(-1,1)$, it holds that
$\chi\geq\frac{1}{p}\geq\frac{(1+p)^{2}}{4p}>\frac{(1-\lambda_{n})(1+p)^{2}}{8p}\Longrightarrow\nu_{i}\geq 1-\frac{1}{2\chi}(1-\lambda_{n})>\frac{(1-p)^{2}}{(1+p)^{2}}.$
As a result, when $\chi\geq\frac{1}{p}$, we have $\nu_{i}\in(c_{i},1)$, i.e., $\Delta_{i}(\nu_{i},p)<0$. It implies that
$\gamma_{(1,2),i}$ $=\frac{1}{2}\big{[}(1+p)\nu_{i}+(1-p)\big{]}\pm j\frac{1}{2}\sqrt{4\nu_{i}-\big{[}(1+p)\nu_{i}+(1-p)\big{]}^{2}},\text{ and }|\gamma_{(1,2),i}|=\sqrt{\nu_{i}}<1,$
where $j^{2}=-1$. Since $\gamma_{1,i}\neq\gamma_{2,i}$, there exists a invertible $Q^{\mathrm{s}}_{2,i}$ such that $H_{i}=Q^{\mathrm{s}}_{2,i}\Gamma_{i}(Q^{\mathrm{s}}_{2,i})^{-1}$, where $\Gamma_{i}=\mathrm{diag}\{\gamma_{1,i},\gamma_{2,i}\}$. Using [5, Appendix B.2] and letting $r=\sqrt{1-\nu_{i}}$, we have
$Q^{\mathrm{s}}_{2,i}=\left[\begin{array}[]{cc}\frac{1}{2}(p-1)\sqrt{1-\nu_{i}}+\frac{1}{2}j\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}&\frac{1}{2}(p-1)\sqrt{1-\nu_{i}}-\frac{1}{2}j\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}\\ p\sqrt{\nicefrac{{1}}{{2\chi}}}&p\sqrt{\nicefrac{{1}}{{2\chi}}}\end{array}\right]$
$(Q^{\mathrm{s}}_{2,i})^{-1}=\frac{\sqrt{2\chi}}{p\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}}\left[\begin{array}[]{cc}-jp\sqrt{\nicefrac{{1}}{{2\chi}}}&\frac{1}{2}\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}+\frac{1}{2}j(p-1)\sqrt{1-\nu_{i}}\\ jp\sqrt{\nicefrac{{1}}{{2\chi}}}&\frac{1}{2}\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}-\frac{1}{2}j(p-1)\sqrt{1-\nu_{i}}\end{array}\right]$
Since the spectral radius of matrix is upper bounded by any of its norm, $10<p_{0}\leq p<$ 1, and $10<\nu_{i}<$ 1, it holds that
$\|Q_{2,i}\|^{2}\leq\|Q_{2,i}Q_{2,i}^{*}\|_{1}\leq 4.$
Following a similar argument for $Q_{2,i}^{-1}$, and using $p^{2}(1+p)^{2}(\nu_{i}-c_{i})=p^{2}(1+p)^{2}(1-\frac{1}{2\chi}(1-\lambda_{i}))-(1-p)^{2}\geq 4p^{3}-\frac{4p^{2}(1-\lambda_{n})}{2\chi}\geq\frac{2p^{2}(1+\lambda_{n})}{\chi}$, we have
$\|(Q_{2,i}^{\mathrm{s}})^{-1}\|^{2}\leq\frac{2\chi}{p^{2}(1+p)^{2}(\nu_{i}-c_{i})}\leq\frac{\chi^{2}}{p^{2}(1+\lambda_{n})}.$
Let ${\bf{Q}}^{\mathrm{s}}=({\bf{Q}}^{\mathrm{s}}_{1})^{\sf T}{\bf{Q}}^{\mathrm{s}}_{2}$ with ${\bf{Q}}^{\mathrm{s}}_{2}=\mathrm{blkdiag}\{Q^{\mathrm{s}}_{2,i}\}_{i=2}^{n}$. We have $({\bf{Q}}^{\mathrm{s}})^{-1}{\bf{H}}{\bf{Q}}^{\mathrm{s}}={\bf{\Gamma}}$, where ${\bf{\Gamma}}=\mathrm{blkdiag}\{\Gamma_{i}\}_{i=2}^{n}$, i.e., there exists an invertible matrix ${\bf{Q}}^{\mathrm{s}}$ such that ${\bf{H}}^{\mathrm{s}}={\bf{Q}}^{\mathrm{s}}{\bf{\Gamma}}({\bf{Q}}^{\mathrm{s}})^{-1}$, and
$\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1.$
Moreover, we have $\|{\bf{Q}}^{\mathrm{s}}\|^{2}\|({\bf{Q}}^{\mathrm{s}})^{-1}\|^{2}\leq\frac{8\chi^{2}}{p^{2}(1+\lambda_{n})}$. We thus complete the proof. $\Box$
XIII Proof of Lemma 6
Proof Proof of (46). It follows from (60) and $0<\alpha L\leq\frac{1}{2}$ that
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{e}}}^{t+1}\big{\|}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq(1-\mu\alpha)\|\bar{{\bf{e}}}^{t}\|^{2}+\Big{(}\frac{\alpha L}{n}+\frac{2\alpha^{2}L^{2}}{n}\Big{)}\|{\bf{X}}^{t}-{\bf{1}}\bar{{\bf{x}}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}-2\alpha(1-2\alpha L)(f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star}))$
$\leq(1-\mu\alpha)\|\bar{{\bf{e}}}^{t}\|^{2}+\frac{2\alpha L}{n}\|{\bf{X}}^{t}-{\bf{1}}\bar{{\bf{x}}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}.$
[…]
$\|\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}=\|\hat{{\bf{P}}}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}=\|({\bf{I}}-\frac{1}{n}{\bf{11}}^{\sf T})\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}=\|{\bf{X}}^{t}-{\bf{1}}\bar{{\bf{x}}}^{t}\|_{\mathrm{F}}^{2}.$
On the other hand, $\|\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}=\|\upsilon^{-1}{\bf{Q}}^{\mathrm{s}}\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}-\|\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\|_{\mathrm{F}}^{2}$. It holds that
$\|{\bf{X}}^{t}-{\bf{1}}\bar{{\bf{x}}}^{t}\|_{\mathrm{F}}^{2}\leq\|\upsilon^{-1}{\bf{Q}}^{\mathrm{s}}\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\leq\upsilon^{-2}\|{\bf{Q}}^{\mathrm{s}}\|^{2}\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}.$
Therefore, we (46) follows.
Proof of (47). Taking conditioned expectation with respect to $\mathcal{F}^{t}$, it follows from (45f) that
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$ $=\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|\mathbb{F}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+2\mathbb{E}\!\left[\langle\mathbb{G}^{t}_{\mathrm{s}},\mathbb{F}^{t}_{\mathrm{s}}\rangle\;|\;\mathcal{F}^{t}\right]$
$=\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|\mathbb{F}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$
$=\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[\|\upsilon p({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right].$
Since ${\bf{E}}^{t}=\frac{(\theta_{t}-1)}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, $\mathop{\rm Prob}(\theta_{t}=1)=p$, and $\mathop{\rm Prob}(\theta_{t}=0)=1-p$, we have
$\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]+\mathbb{E}\!\left[\|\upsilon p({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]$
$=\frac{1-p}{4\chi^{2}}\Big{(}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}+\|\upsilon p({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\Big{)}$
$=\frac{1-p}{4\chi^{2}}\Big{(}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}(\hat{{\bf{Z}}}^{t}-{\bf{X}}^{\star})\|_{\mathrm{F}}^{2}+\|\upsilon p({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}(\hat{{\bf{Z}}}^{t}-{\bf{X}}^{\star})\|_{\mathrm{F}}^{2}\Big{)}$
$\leq\frac{(1-p)(2+p^{2})}{2\chi^{2}}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
Hence, it gives that
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]\leq\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\frac{(1-p)(2+p^{2})}{2\chi^{2}}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
Taking conditioned expectation with respect to $\mathcal{G}^{t}\subset\mathcal{F}^{t}$, and using the unbiasedness of ${\bf{G}}^{t}$, we have
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{(1-p)(2+p^{2})}{2\chi^{2}}\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right].$ (81)
Let $\upsilon=1/\|({\bf{Q}}^{\mathrm{s}})^{-1}\|$. $\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ can be bounded as follows:
$\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=\mathbb{E}\!\left[\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$=\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}+\upsilon^{2}\alpha^{2}\mathbb{E}\!\left[\left\|({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\leq\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}+\frac{(p^{2}+2\chi^{2})n\alpha^{2}\sigma^{2}}{2\chi^{2}}\ .$
The last inequality holds due to $\|\hat{{\bf{\Lambda}}}_{a}\|\leq 1$, $\|\hat{{\bf{\Lambda}}}_{b}\|^{2}\leq 2$, and $\upsilon=\|({\bf{Q}}^{\mathrm{s}})^{-1}\|$. For any vectors ${\bf{a}}$ and ${\bf{b}}$, it holds from Jensen’s inequality that $\|{\bf{a+b}}\|^{2}\leq\frac{1}{\theta}\|{\bf{a}}\|^{2}+\frac{1}{1-\theta}\|{\bf{b}}\|^{2}$ for any $\theta\in(0,1)$. Therefore, letting $\theta=\|{\bf{\Gamma}}\|:=\gamma$, it holds that
$\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}$
$\leq\frac{1}{\gamma}\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}(2\chi^{2}+p^{2})}{2\chi^{2}(1-\gamma)}\|\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})\|_{\mathrm{F}}^{2}$
$\leq\gamma\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}L^{2}(2\chi^{2}+p^{2})}{2\chi^{2}(1-\gamma)}\|{\bf{X}}^{t}-{\bf{X}}^{\star}\|_{\mathrm{F}}^{2}\ .$
Then, we have
$\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq\gamma\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}L^{2}(2\chi^{2}+p^{2})}{2\chi^{2}(1-\gamma)}\|\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+\frac{(p^{2}+2\chi^{2})n\alpha^{2}\sigma^{2}}{2\chi^{2}}\ .$ (82)
In addition, we bound $\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ as follows:
$\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}(\widetilde{{\bf{X}}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})-{\bf{W}}_{b}\widetilde{{\bf{U}}}^{t})\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$=\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}(\widetilde{{\bf{X}}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))-{\bf{W}}_{b}\widetilde{{\bf{U}}}^{t})\|_{\mathrm{F}}^{2}+\mathbb{E}\!\left[\alpha^{2}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\leq 3\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+3\alpha^{2}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\|_{\mathrm{F}}^{2}+3\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\widetilde{{\bf{U}}}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}$
$\leq 3\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+6\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\|_{\mathrm{F}}^{2}+3\alpha^{2}L^{2}\|{\bf{X}}^{t}-{\bf{X}}^{\star}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}$
$\leq 6\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+3\alpha^{2}L^{2}\|\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}.$ (83)
Therefore, substituting (82) and (XIII) into (81), we can conclude (47). $\Box$
XIV Proofs of Theorem 6
Proof From [49, eq. (27)], we have
$\mathbb{E}\!\left[\big{\|}{\bf{X}}^{t+1}-{\bf{X}}^{\star}\big{\|}_{\mathrm{F}}^{2}\;|\;\mathcal{F}_{t}\right]+\frac{2\chi\alpha^{2}}{p^{2}}\mathbb{E}\!\left[\big{\|}{\bf{U}}^{t+1}-{\bf{U}}^{\star}\big{\|}_{\mathrm{F}}^{2}\;|\;\mathcal{F}_{t}\right]$
$\leq\big{\|}\tilde{{\bf{V}}}^{t}-{\bf{V}}^{\star}\big{\|}_{\mathrm{F}}^{2}+\alpha^{2}\Big{(}\frac{2\chi}{p^{2}}-(1-\lambda_{2})\Big{)}\big{\|}{\bf{U}}^{t}-{\bf{U}}^{\star}\big{\|}_{\mathrm{F}}^{2}\ .$ (84)
Then, recalling the definition of $\tilde{{\bf{V}}}^{t}$ and ${\bf{V}}^{\star}$, it gives that
$\big{\|}\tilde{{\bf{V}}}^{t}-{\bf{V}}^{\star}\big{\|}_{\mathrm{F}}^{2}$ $=\big{\|}({\bf{X}}^{t}-\alpha\nabla F({\bf{X}}^{t}))-({\bf{X}}^{\star}-\alpha\nabla F({\bf{X}}^{\star}))+(\alpha\nabla F({\bf{X}}^{t})-\alpha{\bf{G}}^{t})\big{\|}_{\mathrm{F}}^{2}$
$=\big{\|}({\bf{X}}^{t}-\alpha\nabla F({\bf{X}}^{t}))-({\bf{X}}^{\star}-\alpha\nabla F({\bf{X}}^{\star}))\big{\|}_{\mathrm{F}}^{2}+\big{\|}\alpha\nabla F({\bf{X}}^{t})-\alpha{\bf{G}}^{t}\big{\|}_{\mathrm{F}}^{2}$
$\quad+2\big{\langle}({\bf{X}}^{t}-\alpha\nabla F({\bf{X}}^{t}))-({\bf{X}}^{\star}-\alpha\nabla F({\bf{X}}^{\star})),\alpha\nabla F({\bf{X}}^{t})-\alpha{\bf{G}}^{t}\big{\rangle}.$
Taking conditioned expectation with respect to $\mathcal{G}^{t}\subset\mathcal{F}^{t}$, and using the unbiasedness of ${\bf{G}}^{t}$, we have
$\mathbb{E}\!\left[\big{\|}\tilde{{\bf{V}}}^{t}-{\bf{V}}^{\star}\big{\|}_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\big{\|}({\bf{X}}^{t}-\alpha\nabla F({\bf{X}}^{t}))-({\bf{X}}^{\star}-\alpha\nabla F({\bf{X}}^{\star}))\big{\|}_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}.$ (85)
By [49, Lemma 1], it gives that when $0<\alpha<2/L$ and $\mu>0$
$\big{\|}({\bf{X}}^{t}-\alpha\nabla F({\bf{X}}^{t}))-({\bf{X}}^{\star}-\alpha\nabla F({\bf{X}}^{\star}))\big{\|}_{\mathrm{F}}^{2}\leq\max\{(1-\alpha\mu)^{2},(\alpha L-1)^{2}\}\big{\|}{\bf{X}}^{t}-{\bf{X}}^{\star}\big{\|}_{\mathrm{F}}^{2},$ (86)
and $\max\{(1-\alpha\mu)^{2},(\alpha L-1)^{2}\}\in(0,1)$. Combining with (85), it gives that
$\mathbb{E}\!\left[\big{\|}\tilde{{\bf{V}}}^{t}-{\bf{V}}^{\star}\big{\|}_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq\max\{(1-\alpha\mu)^{2},(\alpha L-1)^{2}\}\big{\|}{\bf{X}}^{t}-{\bf{X}}^{\star}\big{\|}_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}.$ (87)
Then, it follows from (XIV) and (87) that
$\mathbb{E}\!\left[\big{\|}{\bf{X}}^{t+1}-{\bf{X}}^{\star}\big{\|}_{\mathrm{F}}^{2}\right]+\frac{2\chi\alpha^{2}}{p^{2}}\mathbb{E}\!\left[\big{\|}{\bf{U}}^{t+1}-{\bf{U}}^{\star}\big{\|}_{\mathrm{F}}^{2}\right]$
$\leq\max\{(1-\alpha\mu)^{2},(\alpha L-1)^{2}\}\big{\|}{\bf{X}}^{t}-{\bf{X}}^{\star}\big{\|}_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}+(\frac{2\chi\alpha^{2}}{p^{2}}-(1-\lambda_{2})\alpha^{2})\big{\|}{\bf{U}}^{t}-{\bf{U}}^{\star}\big{\|}_{\mathrm{F}}^{2}$
$\leq\max\{(1-\mu\alpha)^{2},(\alpha L-1)^{2},1-\frac{(1-\lambda_{2})p^{2}}{2\chi}\}\Big{(}\|{\bf{X}}^{t}-{\bf{X}}^{\star}\|_{\mathrm{F}}^{2}+\frac{2\chi\alpha^{2}}{p^{2}}\|{\bf{U}}^{t}-{\bf{U}}^{\star}\|_{\mathrm{F}}^{2}\Big{)}+n\alpha^{2}\sigma^{2}$
$=\underbrace{\max\{1-(2\mu\alpha-\mu^{2}\alpha^{2}),1-(2\alpha L-\alpha^{2}L^{2}),1-\frac{(1-\lambda_{2})p^{2}}{2\chi}\}}_{:=\zeta}\Big{(}\|{\bf{X}}^{t}-{\bf{X}}^{\star}\|_{\mathrm{F}}^{2}+\frac{2\chi\alpha^{2}}{p^{2}}\|{\bf{U}}^{t}-{\bf{U}}^{\star}\|_{\mathrm{F}}^{2}\Big{)}+n\alpha^{2}\sigma^{2}.$
Since $0<\alpha<\frac{2}{L}$, $10<\frac{1-\lambda_{2}}{2}<$ 1 and $0<p^{2}\leq 1$, we have $10<\zeta<$ 1. It follows from $\Psi^{t}=\|{\bf{X}}^{t}-{\bf{X}}^{\star}\|_{\mathrm{F}}^{2}+\frac{2\chi\alpha^{2}}{p^{2}}\|{\bf{U}}^{t}-{\bf{U}}^{\star}\|_{\mathrm{F}}^{2}$ that
$\mathbb{E}\!\left[\Psi^{t+1}\right]\leq\zeta\Psi^{t}+n\alpha^{2}\sigma^{2}.$
Taking full expectation, and unrolling the recurrence, we have
$\mathbb{E}\!\left[\Psi^{T}\right]\leq\zeta^{T}\Psi^{0}+\frac{n\alpha^{2}\sigma^{2}}{1-\zeta}.$ (88)
Note that
$1-\frac{p^{2}}{2\chi\kappa_{w}}=1-\frac{p^{2}(1-\lambda_{2})}{2\chi}\leq\sqrt{1-\frac{p^{2}(1-\lambda_{2})}{2\chi}}<1,\text{ and }\gamma=\sqrt{1-\frac{1-\lambda_{2}}{2\chi}}.$
Since $\tilde{\gamma}_{\mathrm{s}}=\gamma+\frac{3(1-p)(2+p^{2})}{\chi^{2}}=\sqrt{1-\frac{1-\lambda_{2}}{2\chi}}+\frac{3(1-p)(2+p^{2})}{\chi^{2}}$, we have
$\chi\geq\frac{36}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}_{\mathrm{s}}\leq\sqrt{1-\frac{p^{2}}{2\chi\kappa_{w}}}=\sqrt{1-\frac{p^{2}(1-\lambda_{2})}{2\chi}}<1.$
From (88), we have $\mathbb{E}[\|\widetilde{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}]\leq\mathbb{E}[\Psi^{t}]\leq\zeta^{t}\Psi^{0}+\frac{n\alpha^{2}\sigma^{2}}{1-\zeta}$. Substituting it to (47), we get
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\right]\leq\tilde{\gamma}_{\mathrm{s}}\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\right]+F_{1}\zeta^{t}+F_{2},$ (89)
where $F_{1}=D_{1}\Psi^{0}$ and $F_{2}=\frac{D_{1}n\alpha^{2}\sigma^{2}}{1-\zeta}+D_{2}n\alpha^{2}\sigma^{2}$. Unrolling the recurrence (89), we have
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\right]$ $\leq\tilde{\gamma}_{\mathrm{s}}\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\right]+F_{1}\zeta^{t}+F_{2}$
$\leq\tilde{\gamma}_{\mathrm{s}}^{t+1}{\|\mathcal{E}_{\mathrm{s}}^{0}\|_{\mathrm{F}}^{2}}+F_{1}\sum_{j=0}^{t}\tilde{\gamma}_{\mathrm{s}}^{j}\zeta^{t-j}+F_{2}\sum_{j=0}^{t}\zeta^{j}$
$=\tilde{\gamma}_{\mathrm{s}}^{t+1}{\|\mathcal{E}_{\mathrm{s}}^{0}\|_{\mathrm{F}}^{2}}+F_{1}\frac{\zeta^{t+1}-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{\zeta-\tilde{\gamma}_{\mathrm{s}}}+F_{2}\frac{1-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{1-\tilde{\gamma}_{\mathrm{s}}}$
$=\tilde{\gamma}_{\mathrm{s}}^{t+1}\Big{(}{\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{0}\|_{\mathrm{F}}^{2}+\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{0}\|_{\mathrm{F}}^{2}}\Big{)}+F_{1}\frac{\zeta^{t+1}-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{\zeta-\tilde{\gamma}_{\mathrm{s}}}+F_{2}\frac{1-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{1-\tilde{\gamma}_{\mathrm{s}}}.$ (90)
Since ${\bf{X}}^{0}=[{\bf{x}}^{0},\cdots,{\bf{x}}^{0}]^{\sf T}$ and ${\bf{U}}^{0}=0$, we have
$\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{0}\|_{\mathrm{F}}^{2}+\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{0}\|_{\mathrm{F}}^{2}\leq\alpha^{2}\|\hat{{\bf{P}}}^{\sf T}{\bf{U}}^{\star}\|_{\mathrm{F}}^{2}.$
Multiplying (44a) by $\hat{{\bf{P}}}^{\sf T}$ and using (27), we have
$0=\alpha\hat{{\bf{P}}}^{\sf T}\nabla F({\bf{X}}^{\star})+\alpha\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{U}}^{\star}.$
Then, it holds that
$\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{0}\|_{\mathrm{F}}^{2}+\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{0}\|_{\mathrm{F}}^{2}\leq\alpha^{2}\|\hat{{\bf{P}}}^{\sf T}{\bf{U}}^{\star}\|_{\mathrm{F}}^{2}\leq\frac{\alpha^{2}}{1-\lambda_{2}}\|\nabla F({\bf{X}}^{\star})\|_{\mathrm{F}}^{2}.$ (91)
Combining (XIV) and (91), and using $11-\tilde{\gamma}_{\mathrm{s}}^{t+1}<$ 1, it gives that
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\right]\leq\tilde{\gamma}_{\mathrm{s}}^{t+1}\frac{\alpha^{2}}{1-\lambda_{2}}\|\nabla F({\bf{X}}^{\star})\|_{\mathrm{F}}^{2}+F_{1}\frac{\zeta^{t+1}-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{\zeta-\tilde{\gamma}_{\mathrm{s}}}+\frac{F_{2}}{1-\tilde{\gamma}_{\mathrm{s}}}.$ (92)
Note that
$\left\{\begin{array}[]{cc}\frac{\zeta^{t+1}-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{\zeta-\tilde{\gamma}_{\mathrm{s}}}\leq\frac{\zeta^{t+1}}{\zeta-\tilde{\gamma}_{\mathrm{s}}},&\zeta>\tilde{\gamma}_{\mathrm{s}};\\ \frac{\zeta^{t+1}-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{\zeta-\tilde{\gamma}_{\mathrm{s}}}\leq\frac{\tilde{\gamma}_{\mathrm{s}}^{t+1}}{\tilde{\gamma}_{\mathrm{s}}-\zeta},&\zeta<\tilde{\gamma}_{\mathrm{s}}.\end{array}\right.$
We have $\frac{\zeta^{t+1}-\tilde{\gamma}_{\mathrm{s}}^{t+1}}{\zeta-\tilde{\gamma}_{\mathrm{s}}}\leq\frac{\tilde{\zeta}_{0}^{t+1}}{|\zeta-\tilde{\gamma}_{\mathrm{s}}|}$, where $\tilde{\zeta}_{0}=\max\{\zeta,\tilde{\gamma}_{\mathrm{s}},1-\mu\alpha\}=\max\{1-\alpha\mu,\sqrt{1-\frac{(1-\lambda_{2})p^{2}}{2\chi}}\}$. Substituting (92) into (46), taking full expectation, and unrolling the recurrence, we have
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{e}}}^{t+1}\big{\|}^{2}\right]$ $\leq(1-\mu\alpha)\|\bar{{\bf{e}}}^{t}\|^{2}+\frac{2\alpha L\vartheta_{\mathrm{s}}}{n}\|\mathcal{E}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\sigma^{2}}{n}$
$\leq(1-\mu\alpha)\|\bar{{\bf{e}}}^{t}\|^{2}+\frac{2\alpha L\vartheta_{\mathrm{s}}}{n}\Big{(}\tilde{\gamma}_{\mathrm{s}}^{t}\frac{\alpha^{2}}{1-\lambda_{2}}\|\nabla F({\bf{X}}^{\star})\|_{\mathrm{F}}^{2}+F_{1}\frac{\tilde{\zeta}_{0}^{t}}{|\zeta-\tilde{\gamma}_{\mathrm{s}}|}+\frac{F_{2}}{1-\tilde{\gamma}_{\mathrm{s}}}\Big{)}+\frac{\alpha^{2}\sigma^{2}}{n}$
$\leq(1-\mu\alpha)\|\bar{{\bf{e}}}^{t}\|^{2}+\frac{2\alpha L\vartheta_{\mathrm{s}}(\frac{\alpha^{2}}{1-\lambda_{2}}\|\nabla F({\bf{X}}^{\star})\|_{\mathrm{F}}^{2}+\nicefrac{{F_{1}}}{{|\zeta-\tilde{\gamma}_{\mathrm{s}}|}})}{n}\tilde{\zeta}_{0}^{t}+\frac{2\alpha L\vartheta_{\mathrm{s}}F_{2}}{n(1-\tilde{\gamma}_{\mathrm{s}})}+\frac{\alpha^{2}\sigma^{2}}{n}$
$\leq\tilde{\zeta}_{0}^{t}a_{0}+\frac{2LF_{2}\vartheta_{\mathrm{s}}}{n\mu(1-\tilde{\gamma}_{\mathrm{s}})}+\frac{\alpha\sigma^{2}}{n\mu}.$
Note that $\chi\geq\frac{72(1-p)}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}_{\mathrm{s}}\leq\frac{1+\gamma}{2}<1$. We have $\frac{1}{1-\tilde{\gamma}_{\mathrm{s}}}\leq\frac{8\chi}{1-\lambda_{2}}$. Since $\vartheta_{\mathrm{s}}=\|{\bf{Q}}^{\mathrm{s}}\|^{2}\|({\bf{Q}}^{\mathrm{s}})^{-1}\|^{2}\leq\frac{8\chi^{2}}{p^{2}(1+\lambda_{n})}$ and $F_{2}=\frac{D_{1}n\alpha^{2}\sigma^{2}}{1-\zeta}+D_{2}n\alpha^{2}\sigma^{2}$, where
$D_{1}=\frac{\alpha^{2}L^{2}(2+p^{2})}{2(1-\gamma)}+\frac{3\alpha^{2}L^{2}(1-p)(2+p^{2})}{2},\ D_{2}=\frac{(2-p)(2+p^{2})}{2},$
we have
$\frac{2\alpha L\vartheta_{\mathrm{s}}F_{2}}{n(1-\tilde{\gamma}_{\mathrm{s}})}\leq\mathcal{O}\left(\frac{\alpha^{4}\sigma^{2}L^{3}\chi^{4}}{\mu p^{2}(1-\lambda_{2})^{2}(1-\zeta)}+\frac{\alpha^{2}\sigma^{2}L\chi^{3}}{\mu p^{2}(1-\lambda_{2})}\right).$
The linear speedup result (49) is thus proved. $\Box$